### v1

Abstract
The celebrated conjecture by Payne, Pólya and Weinberger (1956) states that for the fixed membrane problem, the ratio of the first two eigenvalues, $\lambda_{2}/\lambda_{1}$, is maximized by a disk. A more general dimensional version of this conjecture was later resolved by Ashbaugh and Benguria in the 1990s. For the Robin Laplacian, Payne and Schaefer (2001) formulated an analogous conjecture, positing that the ratio $\mu_{2}/\mu_{1}$ is also maximized by a disk for a range of the boundary parameter $\sigma$. This was later restated by Henrot in 2003. In this work, we affirm this conjecture for all dimensions $N\geq 2$ and for all $\sigma$ greater than or equal to a critical positive value $\sigma_{*}$ which may depend on $\Omega$. Furthermore, we prove that the maximum value of $\mu_{2}/\mu_{1}$ is strictly decreasing in $\sigma$ over the entire interval $(0,+\infty)$. Our result provides a positive answer to a variant of Yau’s Problem 77: by measuring the ratio of the first two eigenfrequencies, one can determine whether an elastically supported drum is circular.
Keywords: PPW type conjecture; Payne-Schaefer conjecture; Spectral gap; Robin problem; Bessel function
AMS Subjection Classification(2020): 15A42; 33C10; 35J25; 35P15
1 Introduction
Let $\lambda_{n}$ with $n\in\mathbb{N}$ denote the eigenvalues (counting multiplicities and arranged in nondecreasing order) of the Dirichlet eigenvalue problem
$\left\{\begin{array}[]{ll}-\Delta u=\lambda u&\text{in}\,\,\,\Omega,\\ u=0&\text{on}\,\,\,\partial\Omega,\end{array}\right$.
where $\Omega\subseteq\mathbb{R}^{N}$ with $N\geq 2$ is a bounded domain (connected open set).
In 1956, Payne, Pólya and Weinberger [42] established a foundational inequality for the eigenvalues of the Dirichlet Laplacian in two dimensions, proving that for all $n\in\mathbb{N}$,
$\lambda_{n+1}-\lambda_{n}\leq\frac{2}{n}\sum_{i=1}^{n}\lambda_{i},\,\,\,n\in\mathbb{N}.$
This result was later extended to arbitrary dimensions $N\geq 2$ by Hile and Protter [30], who derived the stronger inequality
$\sum_{i=1}^{n}\frac{\lambda_{i}}{\lambda_{n+1}-\lambda_{i}}\geq\frac{nN}{4},\,\,\,n\in\mathbb{N}.$
A further refinement was achieved by Yang [47] with the following, sharper bound
$\sum_{i=1}^{n}\left(\lambda_{n+1}-\lambda_{i}\right)\left(\lambda_{n+1}-\left(1+\frac{4}{N}\right)\lambda_{i}\right)\leq 0,\,\,\,n\in\mathbb{N}.$
In the particular case $N=2$ and $n=1$, any of these inequalities yields the universal bound
$\frac{\lambda_{2}}{\lambda_{1}}\leq 3.$
Payne, Pólya and Weinberger conjectured that this upper bound is not sharp, and that the maximum value of $\lambda_{2}/\lambda_{1}$ is actually attained when $\Omega$ is a disk. This leads to the celebrated PPW conjecture
$\frac{\lambda_{2}}{\lambda_{1}}\leq\frac{\lambda_{2}}{\lambda_{1}}\Big|_{\text{disk}}=\frac{j_{1,1}^{2}}{j_{0,1}^{2}}\approx 2.5387,$
where $j_{\nu,k}$ denotes the $k$th positive zero of the Bessel function $J_{\nu}(t)$ (here we use the notions of [1]).
A notable inclusion of the PPW conjecture is found in Yau’s problem list [48, problem 77]. Yau highlighted its profound implication: verifying the conjecture would mean that the first two tones of a drum uniquely characterize its circular shape.
The upper bound of $3$ for this ratio has been progressively improved by several mathematicians. Notably, Brands [11] improved it to $2.686$, de Vries [46] further advanced it to $2.658$, and Chiti [15] ultimately reached $2.586$. In higher dimensions $N\geq 2$, the generalized PPW conjecture states that
$\frac{\lambda_{2}}{\lambda_{1}}\leq\frac{\lambda_{2}}{\lambda_{1}}\Big|_{N-\text{dimensional ball}}=\frac{j_{N/2,1}^{2}}{j_{N/2-1,1}^{2}}.$
This conjecture was conclusively proved in the early 1990s by Ashbaugh and Benguria [6, 7, 8]. For convenience, we will refer to this result as the Ashbaugh-Benguria inequality in the subsequent discussion
Consider the eigenvalue problem for the Laplacian with a Robin boundary condition
$\left\{\begin{array}[]{ll}-\Delta u=\mu u&\text{in}\,\,\,\Omega,\\ \partial_{\texttt{n}}u+\sigma u=0&\text{on}\,\,\,\partial\Omega.\end{array}\right$. (1.1)
Here, $\Omega\subset\mathbb{R}^{N}$ is a bounded domain (connected open set), $\sigma>0$ is a physical constant, and $\mathbf{n}$ is the unit outer normal. We denote by $\mu_{n}$ with $n\in\mathbb{N}$ the $n$th eigenvalue, where the sequence $\mu_{n}$ is arranged in nondecreasing order and accounts for multiplicities.
The Robin condition models a situation where the flux is proportional to the temperature itself [16]. For general theory, we refer to [29, Chapter 4] and [9] and its references; for recent advances, see [4, 19, 20, 21, 22, 23, 33, 34].
In 2001, Payne and Schaefer [39] formulated a conjecture for the two-dimensional elastically supported membrane problem, analogous to the classical PPW conjecture for the fixed membrane problem.
Payne-Schaefer conjecture. The ratio $\mu_{2}/\mu_{1}$ achieves its maximum for the disk for all values of $\sigma$ or for a range of values of $\sigma$.
They proved that $\mu_{2}/\mu_{1}\leq 3$ for all $\sigma\geq\widetilde{\sigma}$, where $\widetilde{\sigma}$ is a positive constant. Subsequently, Henrot [28, Open problem 15] restated this conjecture as the following open problem.
Open problem. For what values of $\sigma$ does the ratio $\mu_{2}/\mu_{1}$ achieve its maximum for the disk?
For general dimensions $N\geq 2$, a natural extension of this problem is to determine whether the ratio $\mu_{2}/\mu_{1}$ is maximized by an $N$-dimensional ball for all $\sigma>0$, or at least for a certain range of $\sigma$. To investigate this, let $B_{R}$ denote the ball of radius $R\in(0,+\infty)$ centered at the origin in $\mathbb{R}^{N}$, and let $\nu=N/2-1$.
In this work, we establish the validity of this conjecture for all dimensions $N\geq 2$ and for all sufficiently large values of the boundary parameter $\sigma$.
Theorem 1.1. Let $\Omega\subset\mathbb{R}^{N}$ ($N\geq 2$) be a given bounded domain whose boundary satisfies an interior sphere condition. There exists a constant $\sigma_{*}(\Omega)>0$ such that for all $\sigma\geq\sigma_{*}$, the ratio of the first two Robin eigenvalues of the Laplacian on $\Omega$ satisfies
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{\mu_{2}}{\mu_{1}}\Big|_{B_{R}}=\frac{k_{\nu+1,1}^{2}}{k_{\nu,1}^{2}},$
where $B_{R}$ is a ball of radius $R$ chosen to have the same volume as $\Omega$ and $k_{\nu+l,m}$ denotes the $m$th positive zero of the function $kJ_{\nu+l+1}(k)-(\sigma+l)J_{\nu+l}(k)$. Furthermore, equality holds if and only if $\Omega$ itself is a ball ($\Omega=B_{R}$).
Theorem 1.1 has a direct physical implication for a drum with an elastically supported boundary: one can indeed determine whether such a drum is circular from its first two eigenfrequencies. Specifically, if the ratio $\mu_{2}/\mu_{1}$ attains the universal upper bound $k_{1,1}^{2}/k_{0,1}^{2}$ which is independent of the shape and size of the drum, then the drum must be circular. Conversely, any deviation of this ratio from the maximum value signifies a non-circular shape. Moreover, the ratio of the first two tones cannot exceed $k_{1,1}^{2}/k_{0,1}^{2}$. This conclusively affirms the claim in Yau’s Problem 77 [48] for the case of an elastic drum.
Our result also provides a nuanced answer to Kac’s famous question [31], “Can one hear the shape of a drum?” While the general answer is negative, our theorem shows that for an elastically supported drum, one can definitively “hear” its circularity from the first two tones.
Let $B$ denote the unit ball in $\mathbb{R}^{N}$ centered at the origin. In our recent work [17], we have established that for $B$, the first and second Robin eigenvalues are given by $\mu_{1}(B)=k_{\nu,1}^{2}$ and $\mu_{2}(B)=k_{\nu+1,1}^{2}$, respectively (a result also confirmed in [22]). For example, in low dimensions, this yields the explicit bounds
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{\mu_{2}}{\mu_{1}}\Big|_{2-\text{dimensional ball}}\approx 3.66726$
and
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{\mu_{2}}{\mu_{1}}\Big|_{3-\text{dimensional ball}}\approx 3.05095.$
when $\sigma=1$.
It is known from [17, 22] that $k_{\nu,1}<j_{\nu,1}$ and $k_{\nu+1,1}<j_{\nu+1,1}$, with the limits satisfying
$\lim_{\sigma\to+\infty}k_{\nu,1}=j_{\nu,1},\quad\lim_{\sigma\to+\infty}k_{\nu+1,1}=j_{\nu+1,1},\quad\text{and thus}\quad\lim_{\sigma\to+\infty}\frac{k_{\nu+1,1}^{2}}{k_{\nu,1}^{2}}=\frac{j_{N/2,1}^{2}}{j_{N/2-1,1}^{2}}.$
Furthermore, as shown in [9, Proposition 4.5], the Robin eigenvalue ratio converges to the Dirichlet ratio in the limit: $\lim_{\sigma\to+\infty}\mu_{2}/\mu_{1}=\lambda_{2}/\lambda_{1}$. Consequently, by taking the limit $\sigma\to+\infty$ in the inequality of Theorem 1.1, we recover the classical Ashbaugh-Benguria inequality for the Dirichlet Laplacian. This demonstrates that our main result constitutes a genuine extension of their celebrated work [6, 7, 8] to the Robin boundary condition.
A counterexample to an Ashbaugh-Benguria type inequality for the Robin problem was recently constructed by Laugesen [34, Conjecture G]. Crucially, the domain in this counterexample violates the interior sphere condition and, even, is non-Lipschitz. Since our Theorem 1.1 explicitly requires the boundary to satisfy an interior sphere condition, our positive result does not conflict with the existence of such a counterexample.
Theorem 1.1 establishes that the maximum value of the ratio $\mu_{2}/\mu_{1}$ is $k_{\nu+1,1}^{2}/k_{\nu,1}^{2}$. An immediate consequence is the following upper bound for the spectral gap
$\mu_{2}-\mu_{1}\leq\left(k_{\nu+1,1}^{2}/k_{\nu,1}^{2}-1\right)\mu_{1}.$
To further understand the behavior of this bound, we analyze its dependence on the boundary parameter $\sigma$. In fact, Laugesen conjectured that $\mu_{2}/\mu_{1}$ is strictly decreasing for all $\sigma>0$ [34, Conjecture B]. As a crucial step towards this conjecture, we prove that the maximal value $k_{\nu+1,1}^{2}/k_{\nu,1}^{2}$ itself is strictly decreasing in $\sigma>0$, thereby verifying Laugesen’s conjecture when the domain is a ball.
Theorem 1.2. The maximum value of $\mu_{2}/\mu_{1}$ is strictly decreasing with respect to $\sigma\in(0,+\infty)$. Furthermore, the function $\sigma\mapsto k_{\nu+1,1}^{2}/k_{\nu,1}^{2}$ is continuously and strictly decreasing on $(0,+\infty)$, and its range spans from $\infty$ (as $\sigma\to 0^{+}$) down to $j_{\nu+1,1}^{2}/j_{\nu,1}^{2}$ (as $\sigma\to+\infty$).
Figure 1: The schematic diagram of $k_{\nu+1,1}^{2}/k_{\nu,1}^{2}$.
Figure 1 provides a schematic illustration of the ratio $k_{\nu+1,1}^{2}/k_{\nu,1}^{2}$ as a function of $\sigma$. A crucial implication of Theorem 1.2 for $N=2$ is the existence of a unique critical value $\sigma^{*}>0$ such that $k_{\nu+1,1}^{2}/k_{\nu,1}^{2}\left(\sigma^{*}\right)=3$. Furthermore, the ratio exceeds $3$ for $\sigma<\sigma^{*}$ and falls below $3$ for $\sigma>\sigma^{*}$. Consequently, when $\Omega$ is a ball, the inequality $\mu_{2}/\mu_{1}\leq 3$ holds if and only if $\sigma\geq\sigma^{*}$. This result precisely identifies the optimal parameter range, thereby refining the earlier estimate in [39].
Our analysis also reveals that $3$ is generally not the sharp upper bound for $\mu_{2}/\mu_{1}$. In fact, for any given $\rho\in\left(j_{\nu+1,1}^{2}/j_{\nu,1}^{2},+\infty\right)$, there exists a unique positive constant $\sigma_{\rho}$ such that
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{k_{\nu+1,1}^{2}}{k_{\nu,1}^{2}}\leq\rho$
for all $\sigma\in\left[\sigma_{\rho},+\infty\right)$. This establishes a tunable upper bound for the eigenvalue ratio, a feature fundamentally distinct from the fixed membrane problem.
Defining the relative gap between $\mu_{1}$ and $\mu_{2}$ as $\left(\mu_{2}-\mu_{1}\right)/\mu_{1}$, Theorem 1.2 implies that its maximum possible value is strictly decreasing in $\sigma\in(0,+\infty)$. Consequently, by adjusting the boundary parameter $\sigma$, one can achieve a desired range for the relative gap in practical applications.
While our overall argument follows the framework established by Ashbaugh and Benguria, the Robin boundary condition necessitates significant modifications to their specific techniques [6, 7, 8]. A key adaptation lies in the choice of parameters for the trial function. We define
$w(r)=\frac{J_{\nu+1}(\beta r)}{J_{\nu}(\alpha r)}\quad\text{for }r\in[0,1],$
where $\alpha=k_{\nu,1}$ and $\beta=k_{\nu+1,1}$ are the first positive zeros of $kJ_{\nu+1}(k)-\sigma J_{\nu}(k)$ and $kJ_{\nu+2}(k)-(\sigma+1)J_{\nu+1}(k)$, respectively. This contrasts with the Dirichlet case [6, 7, 8], where $\alpha=j_{\nu,1}$ and $\beta=j_{\nu+1,1}$ were used. This crucial difference ensures that our function $w(r)$ remains regular at $r=1$, unlike the singular function in the original work. Following [6, 7, 8], we also define
$B(r)=w^{\prime}(r)^{2}+(2\nu+1)\frac{w^{2}(r)}{r^{2}}.$
However, due to our different parameter selection, the properties of $B(r)$ here differ substantially from its counterpart in the Dirichlet setting.
In Section 2, we will prove that the function $w(r)$ is strictly increasing on $\left[0,1\right]$, while $B(r)$ is strictly decreasing on the same interval. Although our parameters differ, the core idea from [8] remains applicable; however, the detailed calculations must be adapted to the new boundary condition. Our results yield a stronger conclusion than that in [8], which only established monotonicity. In fact, applying our argument to their setting would also improve their monotonicity result to strict monotonicity. While this stronger conclusion was not necessary in their context, it is crucial for our analysis. Furthermore, in Section 2, we demonstrate that as $\sigma\rightarrow+\infty$, the positive eigenfunctions corresponding to $\mu_{1}$ converge uniformly to zero on $\partial\Omega$, a property essential for applying the interior sphere condition. This convergence allows us to control the maximum value of these eigenfunctions on the boundary by choosing a sufficiently large boundary parameter $\sigma$.
Another significant difference concerns the application of the Chiti comparison result. To set the stage, let us briefly recall its content. Let $\widetilde{u}_{1}$ be the positive eigenfunction corresponding to $\lambda_{1}$, and define $\widetilde{z}(r)=r^{-\nu}J_{\nu}\left(\sqrt{\lambda_{1}}r\right)$. Normalize $\widetilde{u}_{1}$ such that
$\int_{\Omega}\widetilde{u}_{1}^{2}\,\text{d}x=\int_{B_{1/\gamma}}\widetilde{z}^{2}\,\text{d}x$
where $\gamma=\sqrt{\lambda_{1}}/\alpha$. Chiti [13, 14, 15] proved—for a class of equations more general than the Laplacian—that there exists $s_{1}\in\left(0,\left|B_{1/\gamma}\right|\right)$ such that the decreasing rearrangements $\widetilde{u}_{1}^{*}(s)$ and $\widetilde{z}^{*}(s)$ (defined in Section 3) satisfy
$\left\{\begin{array}[]{ll}\widetilde{z}^{*}(s)\geq\widetilde{u}_{1}^{*}(s)&\text{for}\,\,\,\left[0,s_{1}\right],\\ \widetilde{z}^{*}(s)\leq\widetilde{u}_{1}^{*}(s)&\text{for}\,\,\,\left[s_{1},\left|B_{1/\gamma}\right|\right],\end{array}\right$.
where $|B_{1/\gamma}|=\text{meas}\left(B_{1/\gamma}\right)$. This result was later extended by Ashbaugh and Benguria [7, Theorem A.2] to a broader class of second-order elliptic equations.
The Chiti comparison result plays a crucial role in the proof of the Ashbaugh-Benguria inequality. However, this result and its associated lemma (see [13, 14, 15] or [7, Theorem A.2]) are not directly applicable to the Robin problem. The fundamental obstacle is that, under the Robin boundary condition, the decreasing rearrangements of the eigenfunctions may not vanish at the right endpoint of their domain. Consequently, the functions $\widetilde{z}^{*}$ and $\widetilde{u}_{1}^{*}$ might not intersect at all, violating a key premise of the classical result. Establishing a new Chiti-type comparison result for our setting presents additional challenges. The original proof relies on a preliminary conclusion (which we term the Chiti comparison lemma): if $\widetilde{z}^{*}(0)=\widetilde{u}_{1}^{*}(0)$, then $\widetilde{z}^{*}(s)\leq\widetilde{u}_{1}^{*}(s)$ on $\left[0,\left|B_{1/\gamma}\right|\right]$. A close examination reveals that the Dirichlet boundary condition is decisive in this argument. Specifically, the proof proceeds by assuming the contrary and deriving a contradiction from the existence of a point $s_{0}\in\left(0,\left|B_{1/\gamma}\right|\right)$ and a constant $c\geq 1$ such that $\widetilde{z}^{*}\left(s_{0}\right)=c\widetilde{u}_{1}^{*}\left(s_{0}\right)$ and
$\widetilde{z}^{*}(s)<c\widetilde{u}_{1}^{*}(s)$
for $s\in\left[0,s_{0}\right)$. However, for the Robin problem, since the rearrangements may remain positive, the point $s_{0}$ could coincide with the endpoint $\left|B_{1/\gamma}\right|$, and the contradiction cannot be obtained. Therefore, the Chiti comparison lemma fails in our context, necessitating the development of new comparison principles.
To realize this, we first investigate the continuity of the decreasing rearrangement $u_{1}^{*}(s)$ ($s\in[0,|\Omega|]$) of the first Robin eigenfunction $u_{1}$. Then we establish the following isoperimetric-type inequality
$\left(u_{1}^{*}\right)^{\prime}(s)\leq\mu_{1}N^{-2}C_{N}^{-2/N}s^{\frac{2}{N}-2}\int_{0}^{s}u_{1}^{*}(t)\,\text{d}t,$
where $C_{N}=\pi^{N/2}/\Gamma\left(1+N/2\right)$ is the volume of the $N$-dimensional unit ball. The principal difficulty in deriving this inequality for the Robin problem—and indeed a central challenge in applying symmetric rearrangement methods to it—stems from the behavior of the eigenfunction on the boundary. Unlike the Dirichlet case, $u_{1}$ is generally non-constant on $\partial\Omega$. Consequently, the level set $\left\{x\in\Omega:u_{1}>t\right\}:=U_{t}$ may include portions of $\partial\Omega$ where $u_{1}>t$, which fundamentally alters its perimeter and geometric properties. This complication prevents a direct application of the techniques used for the fixed membrane problem [7, 13, 43], as noted in [2].
To overcome this challenge, we adapt an elegant idea from Alvino et al. [2, 3], which involves decomposing the boundary $\partial U_{t}$ of a level set into its internal part and its intersection with $\partial\Omega$. By selecting a sufficiently large boundary parameter $\sigma$, we can ensure that a specific threshold level exceeds the maximum value of $u_{1}$ on $\partial\Omega$. This allows us to exclude the external part of $\partial U_{t}$, thereby establishing the required isoperimetric-type inequality on a suitable subinterval—a result sufficient for our purposes. Building on this, we then derive a new Chiti-type comparison theorem. All these developments are presented in Section 3. Furthermore, leveraging the Faber-Krahn inequality for the Robin problem (established by Bossel [10] for $N=2$ and Daners [18] for $N\geq 2$), we also prove a key domain comparison result in this section, which serves as another essential prerequisite for the proof of Theorem 1.1.
The proof of Theorem 1.1 is presented in Section 4. Owing to the fundamental differences between our new Chiti-type comparison result and the classical version in [7, Theorem A.2], the structure of our proof diverges significantly from the original approach. In particular, establishing the necessity of the equality condition requires more complex and refined arguments. Finally, the proof of Theorem 1.2 is completed in the last section.
2 Strict monotonicity of $w(r)$ and $B(r)$
Consider the functions defined for
$w(r)=\frac{J_{\nu+1}(\beta r)}{J_{\nu}(\alpha r)}\quad\text{and}\quad B(r)=w^{\prime}(r)^{2}+(2\nu+1)\frac{w^{2}(r)}{r^{2}}.$
The values of $w(r)$ and $B(r)$ at $r=0$ are understood as their limits from within the interval $(0,1)$. These functions possess the following two monotonicity results.
Proposition 2.1. The function $w(r)$ is increasing on the interval $\left[0,1\right]$, while $B(r)$ is decreasing on $\left[0,1\right]$.
For the fixed membrane problem (Dirichlet boundary condition), these monotonicity properties were established in [8]. Although our parameters $\alpha$ and $\beta$ differ from their Dirichlet counterparts $\alpha=j_{\nu,1}$ and $\beta=j_{\nu+1,1}$, we successfully prove that the same monotonicity properties hold for the Robin problem. However, due to the fundamental difference in the parameters, our result cannot be directly deduced from [8] and requires an independent proof.
We now present the proof of these monotonicity properties, which involves several notable modifications to the arguments in [8]. Define the auxiliary function
$q(r)=\frac{rw^{\prime}(r)}{w(r)}.$
Following the derivation in [8], we obtain the expressions
$B(r)=\left[q^{2}+(2\nu+1)\right]\left(w/r\right)^{2}$ (2.1)
and
$B^{\prime}(r)=2\left[qq^{\prime}+(q-1)\left(q^{2}+(2\nu+1)\right)/r\right]\left(w/r\right)^{2}.$ (2.2)
The desired conclusions of Proposition 2.1 will follow directly from (2.1) and (2.2) if we can establish that $0\leq q(r)\leq 1$ and $q^{\prime}(r)\leq 0$ for all $r\in\left[0,1\right]$. For $r\in(0,1]$, as shown in [8], the function $q(r)$ can be expressed as
$q(r)=r\left[\beta\frac{J^{\prime}_{\nu+1}}{J_{\nu+1}}(\beta r)-\alpha\frac{J^{\prime}_{\nu}}{J_{\nu}}(\alpha r)\right]$ (2.3)
and its derivative is given by
$q^{\prime}(r)=\left(\alpha^{2}-\beta^{2}\right)r+(1-q)\left(q+(2\nu+1)\right)/r+2\alpha q\frac{J_{\nu+1}}{J_{\nu}}(\alpha r).$ (2.4)
We begin by analyzing the boundary behavior of $q$ at $r=0$ and $r=1$. Note that $q$ is singular only at the left endpoint, since $J_{\nu}(\alpha)>0$ and $J_{\nu+1}(\beta)>0$. Substituting the derivative formula for Bessel functions (see [36])
$J_{\nu}^{\prime}(z)=-J_{\nu+1}(z)+\frac{\nu}{z}J_{\nu}(z)$ (2.5)
into equation (2.3), we obtain the simplified expression
$q(r)=-\beta r\frac{J_{\nu+2}}{J_{\nu+1}}(\beta r)+\alpha r\frac{J_{\nu+1}}{J_{\nu}}(\alpha r)+1.$ (2.6)
Using the limiting form
$J_{\nu}(z)\sim\left(\frac{z}{2}\right)^{\nu}\frac{1}{\Gamma(\nu+1)},\,\,z\rightarrow 0$ (2.7)
in (2.6), we find that $q(0)=1$, where the value is understood in the limit sense as $r\rightarrow 0^{+}$. Furthermore, it follows from (2.4) that $q^{\prime}(0)=0$.
Combining this with (2.4), we derive that
$q^{\prime}(1)=2\nu+1+\alpha^{2}-\beta^{2}+2\sigma q(1)=2\nu+1+\alpha^{2}-\beta^{2},$
which is different from [8].
Differentiating equation (2.4) at $r=0$ yields
$q^{\prime\prime}(0)=\alpha^{2}/(\nu+1)-\beta^{2}/(\nu+2).$
Recall that $\alpha$ and $\beta$ are defined as the first positive zeros of $kJ_{\nu+1}(k)-\sigma J_{\nu}(k)$ and $kJ_{\nu+2}(k)-(\sigma+1)J_{\nu+1}(k)$, respectively. This definition provides the key boundary relations
$\frac{\alpha J_{\nu+1}(\alpha)}{J_{\nu}(\alpha)}=\sigma$ (2.8)
and
$\frac{\beta J_{\nu+2}(\beta)}{J_{\nu+1}(\beta)}=\sigma+1,$ (2.9)
which play a pivotal role in our subsequent analysis and mark a fundamental difference from [8]. Substituting (2.8) and (2.9) into equation (2.6) at $r=1$, we obtain
$q(1)=-\beta\frac{J_{\nu+2}}{J_{\nu+1}}(\beta)+\alpha\frac{J_{\nu+1}}{J_{\nu}}(\alpha)+1=0.$
Combining this result with equation (2.4), we find the derivative at the right endpoint
$q^{\prime}(1)=2\nu+1+\alpha^{2}-\beta^{2}+2\sigma q(1)=2\nu+1+\alpha^{2}-\beta^{2},$
which again differs from the expression in [8].
We next prove that $q^{\prime}(1)<0$ and $q^{\prime\prime}(0)<0$. The proofs of these inequalities differ significantly from those in [8], as they rely crucially on the specific form of the Robin boundary condition.
Lemma 2.1. For $\nu>-1$, the inequality
$2\nu+1+\alpha^{2}-\beta^{2}<0$
holds, which immediately implies $q^{\prime}(1)<0$.
Proof. Note that $\alpha^{2}$ is the first eigenvalue of the problem
$\left(r^{N-1}u^{\prime}(r)\right)^{\prime}=\mu r^{N-1}u\,\,\,\text{on}\,\,\,(0,1]$
subject to the boundary conditions $u^{\prime}(0)=0$ and $u^{\prime}(1)+\sigma u(1)=0$. The corresponding eigenfunction is $u_{1}=r^{-\nu}J_{\nu}(\alpha r)$ (up to a constant factor). By the Rayleigh-Ritz variational principle, we have the inequality
$\alpha^{2}\leq\frac{\sigma u(1)^{2}+\int_{0}^{1}r^{N-1}u^{\prime 2}\,\text{d}r}{\int_{0}^{1}r^{N-1}u^{2}\,\text{d}r}$ (2.10)
for any nontrivial trial function $u$ satisfying the boundary conditions.
We choose the trial function $u=r^{-\nu}J_{\nu+1}(\beta r)$, which is linearly independent of $u_{1}$. Substituting this into inequality (2.10) and applying integration by parts, we derive the following estimate
$\alpha^{2}<$ $\frac{\sigma J_{\nu+1}^{2}(\beta)+\int_{0}^{1}r^{N-1}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{\prime 2}\,\text{d}r}{\int_{0}^{1}r^{N-1}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{2}\,\text{d}r}$
$=$ $\frac{\sigma J_{\nu+1}^{2}(\beta)+\int_{0}^{1}r^{N-1}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{\prime}\,\text{d}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)}{\int_{0}^{1}r^{N-1}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{2}\,\text{d}r}$
$=$ $\frac{-\int_{0}^{1}r^{-\nu}J_{\nu+1}(\beta r)\left(r^{N-1}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{\prime}\right)^{\prime}\,\text{d}r}{\int_{0}^{1}r^{N-1}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{2}\,\text{d}r}$
$=$ $\frac{\beta^{2}\int_{0}^{1}r^{N-1}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{2}\,\text{d}r-(N-1)\int_{0}^{1}r^{N-3}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{2}\,\text{d}r}{\int_{0}^{1}r^{N-1}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{2}\,\text{d}r}$
$=$ $\beta^{2}-(N-1)\frac{\int_{0}^{1}r^{N-3}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{2}\,\text{d}r}{\int_{0}^{1}r^{N-1}\left(r^{-\nu}J_{\nu+1}(\beta r)\right)^{2}\,\text{d}r}$
$<$ $\beta^{2}-(2\nu+1).$
Here, the second equality holds because $r^{-\nu}J_{\nu+1}(\beta r)$ satisfies the boundary condition. The third equality follows from the differential equation (see [17])
$r^{1-N}\left(r^{N-1}u^{\prime}\right)^{\prime}-(N-1)r^{-2}u+\beta^{2}u=0,$
which is satisfied by $u(r)=r^{-\nu}J_{\nu+1}(\beta r)$. The final inequality is due to the fact that $1/r>r$ for $r\in(0,1)$.∎
Remark 2.1. Lemma 2.1 also yields a lower bound for the gap between the first two eigenvalues when $\Omega=B$, namely,
$\mu_{2}-\mu_{1}>2\nu+1,$
which is a noteworthy result in its own right. Taking the limit $\sigma\rightarrow+\infty$, we further obtain
$\lambda_{2}-\lambda_{1}\geq 2\nu+1.$
for the Dirichlet Laplacian on the ball. It is well known that Yau [48, Theorem 7.2 in Chapter 3] established the estimate
$\lambda_{2}-\lambda_{1}\geq\frac{\pi^{2}}{4d^{2}}$
for bounded smooth convex domains, where $d$ is the diameter. For the unit ball ($d=2$), our bound $2\nu+1$ is clearly sharper than Yau’s $\pi^{2}/16$, since $2\nu+1>\pi^{2}/16$.
Lemma 2.2. For $\nu>-1$, the inequality
$\frac{\alpha^{2}}{\nu+1}-\frac{\beta^{2}}{\nu+2}<0$
holds, which implies $q^{\prime\prime}(0)<0$.
Proof. By the Rayleigh-Ritz variational principle, the eigenvalue $\alpha^{2}$ satisfies
$\alpha^{2}\leq\frac{-\int_{0}^{1}\left(r^{N-1}u^{\prime}\right)^{\prime}u\,\text{d}r}{\int_{0}^{1}r^{N-1}u^{2}\,\text{d}r}$ (2.11)
for any admissible trial function $u$. Equality in (2.11) is attained when $u=r^{-\nu}J_{\nu}(\alpha r)$.
Consider the trial function $u=r^{-\nu-1}J_{\nu+1}(\beta r)$. We claim that $r^{-\nu-1}J_{\nu+1}(\beta r)$ and $r^{-\nu}J_{\nu}(\alpha r)$ are linearly independent on $(0,1]$. Suppose, for contradiction, that there exists a nonzero constant $C$ such that
$\frac{J_{\nu+1}(\beta r)}{rJ_{\nu}(\alpha r)}=C.$
Differentiating both sides with respect to $r$ yields
$0$ $=\left[\frac{1}{r}\frac{J_{\nu+1}(\beta r)}{J_{\nu}(\alpha r)}\right]^{\prime}$
$=\frac{\beta rJ^{\prime}_{\nu+1}(\beta r)J_{\nu}(\alpha r)-J_{\nu+1}(\beta r)J_{\nu}(\alpha r)-\alpha rJ_{\nu+1}(\beta r)J^{\prime}_{\nu}(\alpha r)}{r^{2}J^{2}_{\nu}(\alpha r)}$
$=\frac{-\beta rJ_{\nu+2}(\beta r)J_{\nu}(\alpha r)+\alpha rJ_{\nu+1}(\beta r)J_{\nu+1}(\alpha r)}{r^{2}J_{\nu}^{2}(\alpha r)},$
where the last equality follows from the recurrence relation (2.5). This implies
$\alpha rJ_{\nu+1}(\beta r)J_{\nu+1}(\alpha r)=\beta rJ_{\nu+2}(\beta r)J_{\nu}(\alpha r).$
Evaluating at $r=1$, we obtain
$\alpha\frac{J_{\nu+1}(\alpha)}{J_{\nu}(\alpha)}=\beta\frac{J_{\nu+2}(\beta)}{J_{\nu+1}(\beta)}.$
However, from the boundary conditions (2.8) and (2.9), the left-hand side equals $\sigma$, while the right-hand side equals $\sigma+1$. This leads to the contradiction $\sigma=\sigma+1$, thereby proving the linear independence.
Substituting $u=r^{-\nu-1}J_{\nu+1}(\beta r)$ into inequality (2.11) and recalling that $\nu=N/2-1$, we obtain
$\alpha^{2}<\frac{-\int_{0}^{1}u\left[r^{2\nu+1}\left(r^{-\nu-1}J_{\nu+1}(\beta r)\right)^{\prime}\right]^{\prime}\,\text{d}r}{\int_{0}^{1}ur^{\nu}J_{\nu+1}(\beta r)\,\text{d}r}.$
Applying the derivative formula (2.5) again, we find
$\left[r^{-\nu-1}J_{\nu+1}(\beta r)\right]^{\prime}=-\beta r^{-\nu-1}J_{\nu+2}(\beta r)$
and
$\left[r^{\nu}J_{\nu+2}(\beta r)\right]^{\prime}=2(\nu+1)r^{\nu-1}J_{\nu+2}(\beta r)-\beta r^{\nu}J_{\nu+3}(\beta r).$
Consequently, the numerator simplifies as follows
$\alpha^{2}$ $<\frac{\int_{0}^{1}u\left(\beta r^{\nu}J_{\nu+2}(\beta r)\right)^{\prime}\,\text{d}r}{\int_{0}^{1}ur^{\nu}J_{\nu+1}(\beta r)\,\text{d}r}$
$=\beta^{2}\frac{\int_{0}^{1}u\left[\frac{2}{\beta}(\nu+1)r^{\nu-1}J_{\nu+2}(\beta r)-r^{\nu}J_{\nu+3}(\beta r)\right]\,\text{d}r}{\int_{0}^{1}ur^{\nu}J_{\nu+1}(\beta r)\,\text{d}r}.$
Therefore, to prove the inequality
$\frac{\alpha^{2}}{\beta^{2}}<\frac{\nu+1}{\nu+2},$
it suffices to establish that
$\frac{\int_{0}^{1}u\left[\frac{2}{\beta}(\nu+1)r^{\nu-1}J_{\nu+2}(\beta r)-r^{\nu}J_{\nu+3}(\beta r)\right]\,\text{d}r}{\int_{0}^{1}ur^{\nu}J_{\nu+1}(\beta r)\,\text{d}r}\leq\frac{\nu+1}{\nu+2}.$
Consider the following integral identity
$\int_{0}^{1}u\left[-(\nu+1)r^{\nu}J_{\nu+1}(\beta r)+\frac{2}{\beta}(\nu+1)(\nu+2)r^{\nu-1}J_{\nu+2}(\beta r)-(\nu+2)r^{\nu}J_{\nu+3}(\beta r)\right]\,\text{d}r$
$=$ $\int_{0}^{1}u\left[-2(\nu+1)(\nu+2)r^{\nu}\frac{J_{\nu+2}(\beta r)}{\beta r}-r^{\nu}J_{\nu+3}(\beta r)+\frac{2}{\beta}(\nu+1)(\nu+2)r^{\nu-1}J_{\nu+2}(\beta r)\right]\,\text{d}r$
$=$ $-\int_{0}^{1}ur^{\nu}J_{\nu+3}(\beta r)\,\text{d}r<0,$
where the first equality follows from the Bessel function identity $J_{\nu+1}(x)+J_{\nu+3}(x)=2(\nu+2)J_{\nu+2}(x)/x$, and the final inequality holds because $\beta<j_{\nu+1,1}$ ensures the integrand remains positive. This completes the proof.∎
With Lemmas 2.1 and 2.2 established, the desired conclusions of Proposition 2.1 follow from a straightforward adaptation of the argument presented in [8, Theorem 2.5].
Furthermore, we can establish the following stronger result.
Proposition 2.2. The function $w(r)$ is strictly increasing on $\left[0,1\right]$, while $B(r)$ is strictly decreasing on the same interval.
Proof. We claim that $q^{\prime}(r)<0$ in $(0,1)$. Suppose, for contradiction, that this is not the case. Since we already have $q^{\prime}(r)\leq 0$ on $[0,1]$, the failure of the strict inequality implies the existence of two distinct points $r_{1}$ and $r_{2}$ with $r_{1}<r_{2}$ such that $q\left(r_{1}\right)=q\left(r_{2}\right)$. Combined with the fact that $q^{\prime}(r)\leq 0$, this forces $q$ to be constant on the interval $\left[r_{1},r_{2}\right]$. Consequently, there exist at least three points $r_{1}$, $r_{2}$ and $r_{3}$ with $0<r_{1}<r_{3}<r_{2}$ for which $q^{\prime}\left(r_{1}\right)=q^{\prime}\left(r_{2}\right)=q^{\prime}\left(r_{3}\right)=0$. Now, define the function $F(r,q)=\left(\alpha^{2}-\beta^{2}\right)r+(1-q)(q+(2\nu+1))/r+2\alpha qJ_{\nu+1}(\alpha r)/J_{\nu}(\alpha r)$, so that $q^{\prime}(r)=F(r,q(r))$. For a fixed value of $q$, it is known from [8, Theorem 2.5] that $F$ is strictly convex in $r$. Let $r_{3}=\lambda r_{1}+(1-\lambda)r_{2}$ for some $\lambda\in(0,1)$. By the strict convexity of $F$ and the fact that $q$ takes a common value at $r_{1}$, $r_{2}$ and $r_{3}$, we derive the contradiction
$0=q^{\prime}\left(r_{3}\right)=F\left(r_{3},q\right)<\lambda F\left(r_{1},q\right)+(1-\lambda)F\left(r_{2},q\right)=\lambda q^{\prime}\left(r_{1}\right)+(1-\lambda)q^{\prime}\left(r_{2}\right)=0.$
This contradiction establishes that $q^{\prime}(r)<0$ in $(0,1)$.
Moreover, since $q^{\prime}(r)<0$ and $0<q<1$ in $(0,1)$, using (2.2), we obtain $B^{\prime}(r)<0$ in $(0,1)$. Therefore, $B(r)$ is strictly decreasing on $\left[0,1\right]$. Given that $q(0)=1$ and $q(1)=0$, it follows that $0<q<1$ for all $r\in(0,1)$. This, in turn, implies $w^{\prime}(r)>0$ in $(0,1)$, establishing that $w(r)$ is strictly increasing on $\left[0,1\right]$. Furthermore, since $q^{\prime}(r)<0$ and $0<q<1$ in $(0,1)$, we deduce from formula (2.2) that $B^{\prime}(r)<0$ in $(0,1)$. Consequently, $B(r)$ is strictly decreasing on $\left[0,1\right]$.∎
Remark 2.2. Since $k_{\nu,1}<j_{\nu,1}$ and $k_{\nu+1,1}<j_{\nu+1,1}$, we have $w(r)>0$ for all $r\in(0,1]$. From the definition of $q$, we obtain the relation
$w^{\prime}(r)=\frac{w(r)q(r)}{r}.$
Using the limiting form (2.7) and the fact that $q(0)=1$, we deduce
$\lim_{r\rightarrow 0}w^{\prime}(r)=\frac{\beta^{\nu+1}}{2\nu\alpha^{\nu}}.$
Similarly, we find
$\lim_{r\rightarrow 0}\frac{w}{r}=\frac{\beta^{\nu+1}}{2\nu\alpha^{\nu}}.$
Consequently, the limit of $B(r)$ as $r\rightarrow 0$ is
$\lim_{r\rightarrow 0}B(r)=2(\nu+1)\frac{\beta^{2\nu+2}}{4\nu^{2}\alpha^{2\nu}}>0.$
At $r=1$, since $q(1)=0$, we have $w^{\prime}(1)=0$, and thus $B(1)=(2\nu+1)w^{2}(1)>0$ for $\nu>-1/2$. In particular, for $N\geq 2$, $B(r)$ is positive on $[0,1]$. Schematic diagrams of $w(r)$ and $B(r)$ are provided in Figure 2.
Figure 2: The schematic diagrams of $w(r)$ and $B(r)$.
We conclude this section by analyzing the asymptotic behavior of the first eigenfunction with respect to the boundary parameter $\sigma$. Let $u_{1}$ be any positive eigenfunction corresponding to $\mu_{1}$, and denote $u_{1,pM}=\max_{\partial\Omega}u_{1}$ and $u_{1,M}=\max_{\Omega}u_{1}$. Since $\partial\Omega$ satisfies an interior sphere condition, the Hopf Lemma [25, Lemma 3.4] implies that $u_{1,pM}<u_{1,M}$.
Proposition 2.3. Let $\Omega\subset\mathbb{R}^{N}$ with $N\geq 2$ be a bounded domain (connected open set) whose boundary satisfies an interior sphere condition. Then $\lim_{\sigma\rightarrow+\infty}u_{1,pM}=0$.
Proof. Since $\mu_{1}$ is the first eigenvalue of problem (1.1) with corresponding eigenfunction $u_{1}$, the pair $\left(\mu_{1},u_{1}\right)$ satisfies the boundary value problem
$\left\{\begin{array}[]{ll}-\Delta u_{1}=\mu_{1}u_{1}&\text{in}\,\,\,\Omega,\\ \partial_{\texttt{n}}u_{1}+\sigma u_{1}=0&\text{on}\,\,\,\partial\Omega.\end{array}\right$. (2.12)
Without loss of generality, we normalize $u_{1}$ so that $\int_{\Omega}u_{1}^{2}(x)\,\text{d}x=1$.
Let $v_{1}$ be the normalized eigenfunction corresponding to the first Dirichlet eigenvalue $\lambda_{1}$, so that $\int_{\Omega}v_{1}^{2}(x)\,\text{d}x=1$. The pair $\left(\lambda_{1},v_{1}\right)$ then satisfies the Dirichlet problem
$\left\{\begin{array}[]{ll}-\Delta v_{1}=\lambda_{1}v_{1}&\text{in}\,\,\,\Omega,\\ v_{1}=0&\text{on}\,\,\,\partial\Omega.\end{array}\right$. (2.13)
Multiplying the Robin problem (2.12) by $v_{1}$ and integrating by parts over $\Omega$ yields
$\int_{\Omega}\nabla u_{1}\cdot\nabla v_{1}\,\text{d}x=\mu_{1}\int_{\Omega}u_{1}v_{1}\,\text{d}x.$ (2.14)
Similarly, multiplying the Dirichlet problem (2.13) by $u_{1}$ and integrating by parts gives
$\int_{\partial\Omega}u_{1}\partial_{\texttt{n}}v_{1}\,\text{d}S+\int_{\Omega}\nabla u_{1}\cdot\nabla v_{1}\,\text{d}x=\lambda_{1}\int_{\Omega}u_{1}v_{1}\,\text{d}x.$ (2.15)
Subtracting (2.15) from (2.14) results in
$\left(\mu_{1}-\lambda_{1}\right)\int_{\Omega}u_{1}v_{1}\,\text{d}x$ $=\int_{\partial\Omega}u_{1}\partial_{\texttt{n}}v_{1}\,\text{d}S.$
Taking the limit as $\sigma\rightarrow+\infty$ and noting both the boundedness of $\int_{\Omega}u_{1}v_{1}\,\text{d}x$ and the convergence $\lim_{\sigma\rightarrow+\infty}\mu_{1}=\lambda_{1}$ [9, Proposition 4.5], we find
$\lim_{\sigma\rightarrow+\infty}\int_{\partial\Omega}u_{1}\partial_{\texttt{n}}v_{1}\,\text{d}S=0.$
Since $u_{1}$ is continuous and normalized, it is bounded. Moreover, the interior sphere condition implies $\partial_{\texttt{n}}v_{1}<0$ on $\partial\Omega$. Consequently, we conclude
$\lim_{\sigma\rightarrow+\infty}u_{1}=0,$
which establishes the desired result.∎
3 Chiti type comparison results
We begin by recalling some fundamental definitions and results concerning spherically symmetric rearrangement (also known as Schwarz symmetrization); for comprehensive treatments, we refer to [2, 7, 5, 32, 27, 41, 43, 44] and the references therein.
Let $u$ be a real-valued measurable function defined on $\Omega$. Its decreasing rearrangement is the function
$u^{*}(s)=\inf\{t\geq 0:\mu(t)<s\},$
where
$\mu(t)=\text{meas}\{x\in\Omega:|u(x)|>t\}.$
The function $u^{*}$ is non-increasing on $0,|\Omega|]$ with $|\Omega|=\text{meas}(\Omega)$. The spherically symmetric decreasing rearrangement of $u$ is then defined as
$u^{\star}(x)=u^{*}\left(C_{N}|x|^{N}\right)$
whose domain is the ball $\Omega^{*}$ centered at the origin with $\left|\Omega^{*}\right|=|\Omega|$. Owing to the equimeasurability of $u$, $u^{*}$ and $u^{\star}$, we have the identity
$\int_{\Omega}u^{2}(x)\,\text{d}x=\int_{0}^{M}\left(u^{*}\right)^{2}(s)\,\text{d}s=\int_{\Omega^{*}}\left(u^{\star}\right)^{2}(x)\,\text{d}x,$ (3.1)
which will be utilized in subsequent arguments.
We claim that $u_{1}$ is strictly positive on $\overline{\Omega}$. Suppose, for contradiction, that there exists a point $x_{0}\in\partial\Omega$ such that $u_{1}\left(x_{0}\right)=0$. The Robin boundary condition $\partial_{\texttt{n}}u_{1}+\sigma u_{1}=0$ then implies $\partial_{\texttt{n}}u_{1}\left(x_{0}\right)=0$. However, this contradicts the Hopf Lemma [25, Lemma 3.4], which applies due to the interior sphere condition of $\partial\Omega$. Hence, $u_{1}>0$ on $\overline{\Omega}$. Consequently, the decreasing rearrangement satisfies $u_{1}^{*}(|\Omega|)>0$, in contrast to the Dirichlet case where $u_{1}^{*}(|\Omega|)=0$. This distinction is crucial: for the Robin problem, $u_{1}$ may remain strictly positive on $\partial\Omega$, whereas for the fixed membrane problem, it vanishes identically on the boundary. We remark that the interior sphere condition is used precisely to establish the strict positivity of $u_{1}$ on $\overline{\Omega}$ and the inequality $u_{1,pM}<u_{1,M}$. If $\partial\Omega$ were merely Lipschitz, $u_{1}$ could vanish at some boundary points, as the positivity on $\partial\Omega$ depends sensitively on the boundary regularity.
Since $u_{1}>0$ on $\partial\Omega$ and attains its minimum there, we have $u_{1}>u_{1,m}>0$ in $\Omega$, where $u_{1,m}=\min_{\overline{\Omega}}u_{1}=\min_{\partial\Omega}u_{1}$. For $t\geq u_{1,m}$, define the superlevel set
$U_{t}=\left\{x\in\Omega:u_{1}(x)>t\right\}$
and its volume $\mu(t)=\left|U_{t}\right|$. The function $\mu(t)$ is right-continuous, as $\lim_{\varepsilon\rightarrow 0+}\mu(t+\varepsilon)=\mu(t)$. By Sard’s Theorem, the set of levels $t$ for which $\left\{x\in\Omega:u_{1}(x)=t\right\}$ contains critical points of $u_{1}$ has measure zero. Letting $\partial U_{t}^{int}=\partial U_{t}\cap\Omega$ and $\partial U_{t}^{ext}=\partial U_{t}\cap\partial\Omega$, one sees that $\partial U_{t}=\partial U_{t}^{int}\cup\partial U_{t}^{ext}$ (see [2] for this decompose [^1] We are indebted to Professor Richard Laugesen for providing us with this seminal work, which laid the foundation for our study.). For almost every $t\geq u_{1,m}$, the coarea formula gives
$\mu^{\prime}(t)=\int_{\partial U_{t}^{int}}\frac{1}{\left|\nabla u_{1}\right|}\,H_{N-1}(\text{d}x),$
where $H_{N-1}$ denotes the (N1)(N-1)-dimensional Hausdorff measure. The perimeter of $U_{t}$ is defined as
$P_{u_{1}}(t)=\text{per}\left(U_{t}\right)=\left|\partial U_{t}\right|=\int_{\partial U_{t}}H_{N-1}(\text{d}x).$
Applying the Gauss-Green formula to the Robin eigenvalue equation yields
$\mu_{1}\int_{\left\{x\in\Omega:u_{1}(x)>t\right\}}u_{1}\,\text{d}x=\int_{\left\{x\in\Omega:u_{1}(x)>t\right\}}\left(-\Delta u_{1}\right)\,\text{d}x=\int_{\partial U_{t}}g(x)H_{N-1}(\text{d}x),$
where
$g(x)=\left\{\begin{array}[]{ll}\left|\nabla u_{1}\right|&\text{if}\,\,x\in\partial U_{t}^{int},\\ \sigma u_{1}&\text{if}\,\,x\in\partial U_{t}^{ext}.\end{array}\right$.
Applying the Cauchy-Schwarz inequality in a manner analogous to [2, 43, 45], we obtain that
$P_{u_{1}}^{2}(t)\leq$ $\int_{\partial U_{t}}g(x)H_{N-1}(\text{d}x)\int_{\partial U_{t}}\frac{1}{g(x)}H_{N-1}(\text{d}x)$
$=$ $\mu_{1}\int_{\left\{x\in\Omega:u_{1}(x)>t\right\}}u_{1}\,\text{d}x\left(\int_{\partial U_{t}^{ext}}\frac{1}{\left|\nabla u_{1}\right|}\,H_{N-1}(\text{d}x)+\int_{\partial U_{t}^{int}}\frac{1}{\sigma u_{1}}\,H_{N-1}(\text{d}x)\right)$
$=$ $\mu_{1}\int_{0}^{\mu(t)}u_{1}^{*}(s)\,\text{d}s\left(-\mu^{\prime}(t)+\int_{\partial U_{t}^{ext}}\frac{1}{\sigma u_{1}}\,H_{N-1}(\text{d}x)\right).$
Combining this with the classical isoperimetric inequality $P_{u_{1}}(t)\geq NC_{N}^{1/N}\mu^{1-1/N}(t)$ [2, 12, 37, 38], we derive that
$N^{2}C_{N}^{2/N}\mu^{2-2/N}(t)\leq$ $\mu_{1}\int_{0}^{\mu(t)}u_{1}^{*}(s)\,\text{d}s\left(-\mu^{\prime}(t)+\int_{\partial U_{t}^{ext}}\frac{1}{\sigma u_{1}}\,H_{N-1}(\text{d}x)\right).$ (3.2)
This establishes the following key result.
Lemma 3.1. The function $u_{1}^{*}(s)$ is positive and continuous on $0,|\Omega|]$. Moreover, one has that
$\left(u_{1}^{*}\right)^{\prime}(s)\leq\mu_{1}N^{-2}C_{N}^{-2/N}s^{\frac{2}{N}-2}\int_{0}^{s}u_{1}^{*}(t)\,\text{d}t$ (3.3)
for almost every $s\in\left[0,\mu\left(u_{1,pM}\right)\right)$.
Proof. It is clear that $u_{1}^{*}(|\Omega|)=u_{1,m}>0$, which establishes the positivity of $u_{1}^{*}(s)$. By definition, $u_{1}^{*}$ is known to be left-continuous [43]. We now prove that it is in fact continuous on the entire interval. Since $u_{1}$ is continuous, the distribution function $\mu(t)$ is well-defined on $\left[u_{1,m},u_{1,M}\right]$, and can be continuously extended to $\left[0,+\infty\right)$. Being decreasing, bounded, and right-continuous, $\mu(t)$ may have only jump discontinuities. At each such jump point $\mu(t)$, we connect the graph of $\mu(t)$ by including the vertical segment $(\mu(t),\mu(t-)]$, thereby forming a continuous curve $C_{\mu}$. By the definition of $u_{1}^{*}(s)$ [43], its graph over $0,|\Omega|]$ is the reflection of $C_{\mu}$ across the line $s=t$. Since $C_{\mu}$ is continuous, the graph of $u_{1}^{*}(s)$ is also continuous. In particular, $u_{1}^{*}(s)$ is continuous on $(0,|\Omega|]$. To establish continuity at $s=0$, note that the continuity of $u_{1}$ implies $\lim_{t\rightarrow u_{1,M}^{+}}\mu(t)=0$, while $\mu(t)>0$ for all $t<u_{1,M}$. Hence,
$\lim_{s\rightarrow 0^{+}}u_{1}^{*}(s)=u_{1,M}=u_{1}^{*}(0),$
which proves right-continuity at the origin. Therefore, $u_{1}^{*}(s)$ is continuous on $0,|\Omega|]$.
Since $u_{1}^{*}$ remains constant on each connected component of the complement of the range of $\mu(t)$, the desired inequality holds trivially when $\mu(t)<s\leq\mu(t-)$. We therefore focus on the case where $s$ lies within the range of $\mu(t)$, in which $u_{1}^{*}$ and $\mu$ are inverse functions of each other. When $s$ is in the range of $\mu(t)$, we may set $t=\mu^{-1}(s)$. For $s<\mu\left(u_{1,pM}\right)$ and $s$ belongs to the range of $\mu(t)$, it follows that $t>u_{1,pM}$, which implies $\partial U_{t}^{ext}=\emptyset$. Consequently, the boundary integral vanishes in (3.2), and the inequality (3.2) simplifies to
$N^{2}C_{N}^{2/N}\mu^{2-2/N}(t)\leq$ $-\mu_{1}\int_{0}^{\mu(t)}u_{1}^{*}(s)\,\text{d}s\mu^{\prime}(t).$
which implies that
$\left(u_{1}^{*}\right)^{\prime}(s)\leq\mu_{1}N^{-2}C_{N}^{-2/N}s^{\frac{2}{N}-2}\int_{0}^{s}u_{1}^{*}(t)\,\text{d}t.$
This is the desired conclusion.∎
Due to the distinct nature of the Robin boundary condition, the level set boundary generally satisfies $\partial U_{t}\neq\left\{x\in\Omega:u_{1}(x)=t\right\}$. Consequently, inequality (3.3) is established only on a subinterval of $0,|\Omega|]$, in contrast to the Dirichlet case where it holds on the entire interval, as in [7, Inequality A.2]. We note that the derivation of [7, Inequality A.2] follows the method in [43, Section 3], and a related result appears as [13, Lemma 1], which itself relies on [43, Inequality 32]. Furthermore, the continuity of $u_{1}$ directly implies the continuity of its decreasing rearrangement $u_{1}^{*}$ on $0,|\Omega|]$.
Let $R=\alpha/\sqrt{\mu_{1}}$. Since $\alpha^{2}$ is the first eigenvalue of the Robin problem on the unit ball $B$, a standard scaling argument shows that $\mu_{1}$ is the first eigenvalue of the corresponding problem on the ball $B_{R}$
$\left\{\begin{array}[]{ll}-\Delta z=\mu z&\text{in}\,\,\,B_{R},\\ \partial_{\texttt{n}}z+\frac{\sigma}{R}z=0&\text{on}\,\,\,\partial B_{R}\end{array}\right$.
with the corresponding eigenfunction
$z(r)=cr^{-\nu}J_{\nu}\left(\sqrt{\mu_{1}}r\right),$
where $c$ is a nonzero positive constant. Given that $\alpha=k_{\nu,1}$, we have $z(R)>0$, which contrasts with the Dirichlet case where the eigenfunction vanishes on the boundary.
We now establish a fundamental relation between the measures of $B_{R}$ and $\Omega$.
Lemma 3.2. It holds that $\left|B_{R}\right|\leq|\Omega|$, with equality if and only if $\Omega=B_{R}$.
Proof. The first Robin eigenvalue admits the variational characterization
$\mu_{1}(\Omega)=\inf_{u\in H^{1}(\Omega)\setminus\{0\}}\frac{\int_{\Omega}|\nabla u|^{2}\,\text{d}x+\sigma\int_{\partial\Omega}u^{2}\,\text{d}S}{\int_{\Omega}u^{2}\,\text{d}x}.$
By the Faber-Krahn inequality for the Robin Laplacian—established by Bossel [10] for $N=2$ and Daners [18] for $N\geq 2$—we have
$\mu_{1}\left(B_{R}\right)=\mu_{1}(\Omega)\geq\mu_{1}\left(\Omega^{*}\right)$
with equality if and only if $\Omega=\Omega^{*}$.
Now, suppose for contradiction that $\left|B_{R}\right|>|\Omega|$ or $B_{R}\supsetneq\Omega^{*}$. Then by the domain monotonicity of $\mu_{1}$ [40], we would have
$\mu_{1}\left(B_{R}\right)\leq\mu_{1}\left(\Omega^{*}\right).$
It follows that $\mu_{1}\left(B_{R}\right)=\mu_{1}\left(\Omega^{*}\right)$. So one has that $B_{R}=\Omega^{*}$ which is a contradiction. Hence, $B_{R}\subseteq\Omega^{*}$ or $\left|B_{R}\right|\leq\left|\Omega^{*}\right|=|\Omega|$. Furthermore, if $\left|B_{R}\right|=|\Omega|$, then $B_{R}=\Omega^{*}$, and thus
$\mu_{1}(\Omega)=\mu_{1}\left(B_{R}\right)=\mu_{1}\left(\Omega^{*}\right),$
which implies $\Omega=\Omega^{*}=B_{R}$.∎
Domain monotonicity generally fails for Robin eigenvalues. Nevertheless, it holds in certain special configurations. A notable example is the inequality
$\mu_{1}(\Omega,\sigma)\geq\mu_{1}\left(\widetilde{B},\sigma\right),$
where $\widetilde{B}$ is a ball circumscribing $\Omega$. This result was established for dimensions two and three in [40], and the argument readily extends to higher dimensions. Alternative proofs can be found in [24, 35]. As noted by Bucur, Freitas, and Kennedy [9], this monotonicity property can also be derived from the Bossel-Daners inequality [9, Theorem 4.23]. Specifically, using the scaling inequality $\mu_{1}(t\Omega)\leq\mu_{1}(\Omega)$ for all $t\geq 1$ and $\sigma>0$ [9, Inequality 4.10], one obtains
[…]
By Lemma 3.2, $\left|B_{R}\right|=|\Omega|$ if and only if $\Omega=B_{R}$. In this case, the conclusion of Theorem 1.1 holds trivially for all $\sigma>0$. We may therefore assume throughout the remainder of the paper that $\left|B_{R}\right|<|\Omega|$.
Let $z^{*}(s)$ denote the decreasing rearrangement of $z$. It is known from [13, Formula 18] or [7, Formula A.5] that $z^{*}(s)$ satisfies the differential equation
$\left(z^{*}\right)^{\prime}(s)=\mu_{1}N^{-2}C_{N}^{-2/N}s^{2/N-2}\int_{0}^{s}z^{*}(t)\,\text{d}t$ (3.4)
for $s\in\left[0,\left|B_{R}\right|\right]$. Since $z(R)>0$, we have $z^{*}\left(\left|B_{R}\right|\right)>0$, in contrast to the Dirichlet case where the eigenfunction vanishes on the boundary.
We normalize the eigenfunction $z$ so that
$\int_{\Omega}u_{1}^{2}\,\text{d}x=\int_{B_{R}}z^{2}\,\text{d}x.$ (3.5)
As noted in the introduction, the positivity $z^{*}\left(\left|B_{R}\right|\right)>0$ implies that the classical Chiti comparison lemma [13, Lemma 2] may fail for the Robin problem. It is therefore necessary to establish a new relation between $z^{*}(s)$ and $u_{1}^{*}(s)$.
Lemma 3.3. For $\sigma\geq\sigma_{*}$ with some $\sigma_{*}>0$, if $z^{*}(0)<u_{1}^{*}(0)$, then $z^{*}(s)<u_{1}^{*}(s)$ for all $s\in\left[0,\left|B_{R}\right|\right)$.
Proof. Clearly, $z^{*}(s)$ is continuous on $\left[0,\left|B_{R}\right|\right]$, and by Lemmas 3.1 and 3.2, so is $u_{1}^{*}(s)$. The assumption $z^{*}(0)<u_{1}^{*}(0)$ implies that $z^{*}(s)<u_{1}^{*}(s)$ in a neighborhood of $s=0$. Suppose, for contradiction, that the conclusion fails. Then the functions must intersect at least once in $\left[0,\left|B_{R}\right|\right)$; let $s_{1}$ denote the first such intersection point (see Figure 3).
Figure 3: The schematic diagrams of $u_{1}^{*}(s)$ (solid line) and $z^{*}(s)$ (dotted line).
Define a function by
$v(s)=\left\{\begin{array}[]{ll}u_{1}^{*}(s)&\text{for}\,\,\,s\in\left[0,s_{1}\right),\\ z^{*}(s)&\text{for}\,\,\,s\in\left[s_{1},\left|B_{R}\right|\right].\end{array}\right$.
Note that $v(s)$ cannot be an eigenfunction corresponding to $\mu_{1}$. Therefore, the Rayleigh quotient satisfies the strict inequality
$\mu_{1}<\frac{\int_{B_{R}}|\nabla g|^{2}\,\text{d}x+\frac{\sigma}{R}\int_{\partial B_{R}}g^{2}\,\text{d}S}{\int_{B_{R}}g^{2}\,\text{d}x},$ (3.6)
where $g(x)=v\left(C_{N}|x|^{N}\right)$.
The function $v(s)$ is precisely the decreasing rearrangement of $g(x)$. By their equimeasurability, we have
$\int_{B_{R}}g^{2}\,\text{d}x=\int_{0}^{\left|B_{R}\right|}v^{2}(s)\,\text{d}s,$ (3.7)
which can also be derived via a change of variables as in [7, Formula A.9]. Furthermore, from [7, Formula A.10], it follows that
$\int_{B_{R}}|\nabla g|^{2}\,\text{d}x=N^{2}C_{N}^{2/N}\int_{0}^{\left|B_{R}\right|}s^{2-2/N}v^{\prime}(s)^{2}\,\text{d}s.$ (3.8)
In light of equation (3.4) and the definition of $v$, for $s\in\left[s_{1},\left|B_{R}\right|\right]$, we obtain the inequality
$-(z^{*})^{\prime}(s)$ $=\mu_{1}N^{-2}C_{N}^{-2/N}s^{2/N-2}\left(\int_{0}^{s_{1}}z^{*}(t)\,\text{d}t+\int_{s_{1}}^{s}z^{*}(t)\,\text{d}t\right)$
$\leq\mu_{1}N^{-2}C_{N}^{-2/N}s^{2/N-2}\left(\int_{0}^{s_{1}}u_{1}^{*}(t)\,\text{d}t+\int_{s_{1}}^{s}z^{*}(t)\,\text{d}t\right)$
$\leq\mu_{1}N^{-2}C_{N}^{-2/N}s^{2/N-2}\int_{0}^{s}v(t)\,\text{d}t.$
That is to say,
$v^{\prime}(s)\leq\mu_{1}N^{-2}C_{N}^{-2/N}s^{2/N-2}\int_{0}^{s}v(t)\,\text{d}t$ (3.9)
for $s\in\left[s_{1},\left|B_{R}\right|\right]$.
By Proposition 2.3 and the right-continuity of $\mu(t)$, we have
$\lim_{\sigma\rightarrow+\infty}\mu\left(u_{1,pM}\right)=\lim_{\sigma\rightarrow+\infty}\mu\left(u_{1,m}\right)=|\Omega|.$
Since $|\Omega|>\left|B_{R}\right|$, there exists $\sigma_{*}>0$ such that
$\left|B_{R}\right|\leq\mu\left(u_{1,pM}\right)$
for all $\sigma\geq\sigma_{*}$. By Lemma 3.1 and the definition of $v$, we have
$v^{\prime}(s)\leq\mu_{1}N^{-2}C_{N}^{-2/N}s^{2/N-2}\int_{0}^{s}v(t)\,\text{d}t$
for almost every $s\in\left[0,s_{1}\right]$. Combining this with inequality (3.9), we conclude that the same inequality holds for almost every $s\in\left[0,\left|B_{R}\right|\right]$. We need to point out that the constant $\sigma_{*}$ is not universal but depends on the domain [^2] It is our pleasure to thank Professor Richard Laugesen for bringing this to our attention..
Substituting the preceding inequality into (3.8) and integrating by parts, we obtain
$\int_{B_{R}}|\nabla g|^{2}\,\text{d}x\leq$ $-\mu_{1}\int_{0}^{\left|B_{R}\right|}v^{\prime}(s)\int_{0}^{s}v(t)\,\text{d}t\,\text{d}s$
$=$ $-\mu_{1}v\left(\left|B_{R}\right|\right)\int_{0}^{\left|B_{R}\right|}v(s)\,\text{d}s+\mu_{1}\int_{0}^{\left|B_{R}\right|}v^{2}(s)\,\text{d}s.$
Substituting $s=\left|B_{R}\right|$ into (3.4) yields
$\left(z^{*}\right)^{\prime}\left(\left|B_{R}\right|\right)=\mu_{1}N^{-2}C_{N}^{-2/N}\left|B_{R}\right|^{2/N-2}\int_{0}^{\left|B_{R}\right|}z^{*}(s)\,\text{d}s.$
Since $z^{*}(s)\leq v(s)$ one has that
$\left(z^{*}\right)^{\prime}\left(\left|B_{R}\right|\right)\leq\mu_{1}N^{-2}C_{N}^{-2/N}\left|B_{R}\right|^{2/N-2}\int_{0}^{\left|B_{R}\right|}v(s)\,\text{d}s,$
which can also be derived from (3.9). Rearranging this inequality yields
$\int_{0}^{\left|B_{R}\right|}v(s)\,\text{d}s\leq\mu_{1}^{-1}N^{2}C_{N}^{2/N}\left|B_{R}\right|^{2-2/N}\left(z^{*}\right)^{\prime}\left(\left|B_{R}\right|\right).$
So we obtain that
$\int_{B_{R}}|\nabla g|^{2}\,\text{d}x\leq$ $N^{2}C_{N}^{2/N}\left|B_{R}\right|^{2-2/N}\left(z^{*}\right)^{\prime}\left(\left|B_{R}\right|\right)z^{*}\left(\left|B_{R}\right|\right)+\mu_{1}\int_{0}^{\left|B_{R}\right|}v^{2}(s)\,\text{d}s,$
where we have used the fact of $v\left(\left|B_{R}\right|\right)=z^{*}\left(\left|B_{R}\right|\right)$.
Using the relation $z(r)=z(x)=z^{*}(s)$ with $s=C_{N}r^{N}$ and $|x|=r$, we obtain $z^{*}\left(\left|B_{R}\right|\right)=z(R)$ and
$\left(z^{*}\right)^{\prime}\left(\left|B_{R}\right|\right)=N^{-1}C_{N}^{-1}R^{1-N}z^{\prime}(R).$
Substituting into the previous inequality yields
$\int_{B_{R}}|\nabla g|^{2}\,\text{d}x\leq$ $N^{2}C_{N}^{2/N}\left|B_{R}\right|^{2-2/N}N^{-1}C_{N}^{-1}R^{1-N}z^{\prime}(R)z(R)+\mu_{1}\int_{0}^{\left|B_{R}\right|}v^{2}(s)\,\text{d}s$
$=$ $NC_{N}R^{N-1}z^{\prime}(R)z(R)+\mu_{1}\int_{0}^{\left|B_{R}\right|}v^{2}(s)\,\text{d}s.$
Applying the Robin boundary condition $z^{\prime}(R)+\sigma/Rz(R)=0$, we find
$\int_{B_{R}}|\nabla g|^{2}\,\text{d}x\leq$ $-\frac{\sigma}{R}NC_{N}R^{N-1}z^{2}(R)+\mu_{1}\int_{0}^{\left|B_{R}\right|}v^{2}(s)\,\text{d}s$
$=$ $-\frac{\sigma}{R}\int_{\partial B_{R}}z^{2}\,\text{d}S+\mu_{1}\int_{0}^{\left|B_{R}\right|}v^{2}(s)\,\text{d}s.$
According to the relations $z$ and $z^{*}$, $g$ and $v$, respectively, we can derive that $z(R)=g(R)$. Combining this with (3.7) yields that
$\int_{B_{R}}|\nabla g|^{2}\,\text{d}x\leq$ $-\frac{\sigma}{R}NC_{N}R^{N-1}z^{2}(R)+\mu_{1}\int_{0}^{\left|B_{R}\right|}v^{2}(s)\,\text{d}s$
$=$ $-\frac{\sigma}{R}\int_{\partial B_{R}}g^{2}\,\text{d}S+\mu_{1}\int_{B_{R}}g^{2}\,\text{d}x.$
Rearranging terms, we obtain
$\frac{\int_{B_{R}}|\nabla g|^{2}\,\text{d}x+\frac{\sigma}{R}\int_{\partial B_{R}}g^{2}\,\text{d}S}{\int_{B_{R}}g^{2}\,\text{d}x}\leq\mu_{1},$
which contradicts (3.6).∎
By the equimeasurability property (3.1), the normalization condition (3.5) is equivalent to
$\int_{0}^{M}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s.$
Now, suppose for contradiction that $z^{*}(0)<u_{1}^{*}(0)$. Then, applying Lemmas 3.2 and 3.3, we obtain the chain of inequalities
$\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s<\int_{0}^{\left|B_{R}\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s\leq\int_{0}^{M}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s,$ (3.10)
which is a contradiction. Therefore, we must have $z^{*}(0)\geq u_{1}^{*}(0)$. This conclusion aligns with the Dirichlet case, although the proof in the Robin setting is considerably more intricate.
We now turn to the case where $z^{*}(0)=u_{1}^{*}(0)$.
Lemma 3.4. When $\sigma\geq\sigma_{*}$ and $z^{*}(0)=u_{1}^{*}(0)$, the following holds: either $z^{*}(s)\geq u_{1}^{*}(s)$ for all $s\in\left[0,\left|B_{R}\right|\right]$, or there exists some $s_{1}\in\left(0,\left|B_{R}\right|\right)$ such that $z^{*}(s)\geq u_{1}^{*}(s)$ for $s\in\left[0,s_{1}\right]$ and $z^{*}(s)\leq u_{1}^{*}(s)$ for $s\in\left(s_{1},\left|B_{R}\right|\right]$.
Proof. If $z^{*}(s)\geq u_{1}^{*}(s)$ for all $s\in\left[0,\left|B_{R}\right|\right]$, the conclusion is done. Otherwise, there exists some $s_{0}\in\left(0,\left|B_{R}\right|\right]$ such that $z^{*}\left(s_{0}\right)<u_{1}^{*}\left(s_{0}\right)$. By the sign-preserving property of continuous functions, there exists an interval $\left[s_{1},s_{2}\right]$ containing $s_{0}$ on which $z^{*}(s)<u_{1}^{*}(s)$ for $s\in\left(s_{1},s_{2}\right)$ and $z^{*}\left(s_{1}\right)=u_{1}^{*}\left(s_{1}\right)$. Furthermore, we take $\left[s_{1},s_{2}\right]$ to be the first such interval, which implies $z^{*}(s)\geq u_{1}^{*}(s)$ in $\left[0,s_{1}\right]$ (see Figure 4(a)). We claim that $\left[s_{1},s_{2}\right]\neq\left[0,\left|B_{R}\right|\right]$. Indeed, if equality holds, then $z^{*}(s)<u_{1}^{*}(s)$ in $\left(0,\left|B_{R}\right|\right)$. This again implies that
$\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s<\int_{0}^{\left|B_{R}\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s\leq\int_{0}^{M}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s,$
which is a contradiction. Hence, we must have that $s_{1}>0$ or $s_{2}<\left|B_{R}\right|$.
We begin with the case $s_{1}>0$. Here, two subcases arise: (i) $z^{*}(s)\equiv u_{1}^{*}(s)$ in $\left[0,s_{1}\right]$ (see Figure 4(b)); (ii) There exists a subinterval $\left[s_{3},s_{4}\right]\subseteq\left[0,s_{1}\right]$ such that $z^{*}(s)>u_{1}^{*}(s)$ in $\left(s_{3},s_{4}\right)$ with $z^{*}\left(s_{3}\right)=u_{1}^{*}\left(s_{3}\right)$ and $z^{*}\left(s_{4}\right)=u_{1}^{*}\left(s_{4}\right)$ (see Figure 4(c)). Note that $s_{4}\leq s_{1}<\left|B_{R}\right|$ since $s_{1}<s_{0}\leq\left|B_{R}\right|$.
If subcase (ii) occurs, we further distinguish:
  • If $s_{2}=\left|B_{R}\right|$, the desired conclusion follows.
  • If $s_{2}<\left|B_{R}\right|$, we consider:
    • $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},\left|B_{R}\right|\right)$, which also yields the conclusion;
    • Or there exists $s_{5}\in\left(s_{1},\left|B_{R}\right|\right)$ such that $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},s_{5}\right)$ and $z^{*}\left(s_{5}\right)=u_{1}^{*}\left(s_{5}\right)$.
In the last scenario, we define
$v(s)=\left\{\begin{array}[]{ll}u_{1}^{*}(s)&\text{for}\,\,\,s\in\left(s_{1},s_{5}\right),\\ z^{*}(s)&\text{for}\,\,\,s\in\left[0,s_{1}\right]\cup\left[s_{5},\left|B_{R}\right|\right].\end{array}\right$.
Then, reasoning as that of Lemma 3.3 with obvious changes, we again derive the following contradiction
$\mu_{1}<\frac{\int_{B_{R}}|\nabla g|^{2}\,\text{d}x+\frac{\sigma}{R}\int_{\partial B_{R}}g^{2}\,\text{d}S}{\int_{B_{R}}g^{2}\,\text{d}x}\leq\mu_{1},$
where $g(x)=v\left(C_{N}|x|^{N}\right)$.
If subcase (i) occurs, we consider:
  • If $s_{2}=\left|B_{R}\right|$, then $z^{*}(s)\equiv u_{1}^{*}(s)$ in $\left[0,s_{1}\right]$ and $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},\left|B_{R}\right|\right)$. In this case, a contradiction follows as in (3.10).
  • Hence, we must have $s_{2}<\left|B_{R}\right|$. We then have
    • $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},\left|B_{R}\right|\right)$ (see Figure 4(d)), which again leads to a contradiction via (3.10);
    • Or there exists $s_{6}\in\left[s_{2},\left|B_{R}\right|\right)$ such that $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},s_{6}\right)$ and $z^{*}\left(s_{6}\right)=u_{1}^{*}\left(s_{6}\right)$ (see Figure 4(e)).
In the latter situation, we define
$v(s)=\left\{\begin{array}[]{ll}u_{1}^{*}(s)&\text{for}\,\,\,s\in\left(s_{1},s_{6}\right),\\ z^{*}(s)&\text{for}\,\,\,s\in\left[0,s_{1}\right]\cup\left[s_{6},\left|B_{R}\right|\right].\end{array}\right$.
Proceeding as in Lemma 3.3 with obvious changes, we again arrive at a contradiction.
We now consider the case $s_{1}=0$. Here, we must have $s_{2}<\left|B_{R}\right|$. As before, two subcases arise:
  • If $z^{*}(s)<u_{1}^{*}(s)$ in $\left(0,\left|B_{R}\right|\right)$, a contradiction is derived as in inequality (3.10).
  • Or there exists $s_{7}\in\left[s_{2},\left|B_{R}\right|\right)$ such that $z^{*}(s)<u_{1}^{*}(s)$ in $\left(0,s_{7}\right)$ and $z^{*}\left(s_{7}\right)=u_{1}^{*}\left(s_{7}\right)$ (see Figure 4(f)).
In the last subcase, we define
$v(s)=\left\{\begin{array}[]{ll}u_{1}^{*}(s)&\text{for}\,\,\,s\in\left(0,s_{7}\right),\\ z^{*}(s)&\text{for}\,\,\,s\in\left[s_{7},\left|B_{R}\right|\right].\end{array}\right$.
Following the reasoning of Lemma 3.3, we again reach a contradiction.∎
(a) $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},s_{2}\right)$
(b) $z^{*}(s)\equiv u_{1}^{*}(s)$ in $\left[0,s_{1}\right]$
(c) $z^{*}(s)>u_{1}^{*}(s)$ in $\left(s_{3},s_{4}\right)$
(d) $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},\left|B_{R}\right|\right)$
(e) $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},s_{6}\right)$
(f) $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},s_{7}\right)$
Figure 4: The schematic diagrams of $u_{1}^{*}(s)$ (solid line) and $z^{*}(s)$ (dotted line).
Remark 3.1. The proof of Lemma 3.4 reveals that under the second alternative, there exists a subinterval $\left[s_{3},s_{4}\right]\subseteq\left[0,s_{1}\right]$ such that $z^{*}(s)>u_{1}^{*}(s)$ in $\left(s_{3},s_{4}\right)$ with equality at the endpoints, while $z^{*}(s)<u_{1}^{*}(s)$ in $\left(s_{1},\left|B_{R}\right|\right)$.
The conclusion of Lemma 3.4 differs markedly from the Chiti comparison lemma [13, Lemma 2], which asserts that $\widetilde{z}^{*}(s)\leq\widetilde{u}_{1}^{*}(s)$. Our result relies on condition (3.5), under which Chiti’s inequality can in fact be strengthened to an equality, $\widetilde{z}^{*}(s)=\widetilde{u}_{1}^{*}(s)$. In this sense, our conclusion is weaker. This relaxation is fundamentally due to the fact that $z^{*}\left(\left|B_{R}\right|\right)>0$.
We now present a Chiti-type comparison result, whose schematic behavior is illustrated in Figure 5.
Theorem 3.1. Let (3.5) hold. For $\sigma\geq\sigma_{*}$, the functions $u_{1}^{*}(s)$ and $z^{*}(s)$ satisfy either
(i) $z^{*}(s)\geq u_{1}^{*}(s)$ in $\left[0,\left|B_{R}\right|\right]$, or
(ii) there exists a point $s_{1}\in\left(0,\left|B_{R}\right|\right)$ such that $z^{*}\left(s_{1}\right)=u_{1}^{*}\left(s_{1}\right)$ and,
$\left\{\begin{array}[]{ll}z^{*}(s)\geq u_{1}^{*}(s)&\text{for}\,\,\,\left[0,s_{1}\right],\\ z^{*}(s)\leq u_{1}^{*}(s)&\text{for}\,\,\,\left[s_{1},\left|B_{R}\right|\right],\end{array}\right$.
where $z^{*}(s)=z(r)$ with $s=C_{N}|x|^{N}$.
Proof. Using Lemmas 3.2–3.3 we have shown that $z^{*}(0)\geq u_{1}^{*}(0)$. If $z^{*}(0)=u_{1}^{*}(0)$, then by Lemma 3.4, the desired conclusion follows immediately. Therefore we only need to consider the case $z^{*}(0)>u_{1}^{*}(0)$. In this case, if $z^{*}(s)\geq u_{1}^{*}(s)$ in $\left[0,\left|B_{R}\right|\right]$, the result is again obtained. Otherwise, there exists a point $s_{0}\in\left(0,\left|B_{R}\right|\right)$ such that $z^{*}\left(s_{0}\right)<u_{1}^{*}\left(s_{0}\right)$. This further implies that there exists a point $s_{1}\in\left(0,s_{0}\right)$ such that $z^{*}\left(s_{1}\right)=u_{1}^{*}\left(s_{1}\right)$. Without loss of generality, we assume $s_{1}$ is the first positive zero of $z^{*}(s)=u_{1}^{*}(s)$. Then $z^{*}(s)>u_{1}^{*}(s)$ on the interval $\left[0,s_{1}\right)$ and it suffices to show that $z^{*}(s)<u_{1}^{*}(s)$ for all $s\in\left(s_{1},\left|B_{R}\right|\right)$. If not, there exists a point $s_{2}\in\left(s_{1},\left|B_{R}\right|\right)$ such that $z^{*}\left(s_{2}\right)=u_{1}^{*}\left(s_{2}\right)$. Again, assuming $s_{2}$ is the second positive zero of $z^{*}(s)=u_{1}^{*}(s)$, we have $z^{*}(s)<u_{1}^{*}(s)$ for all $s\in\left(s_{1},s_{2}\right)$ with $z^{*}\left(s_{2}\right)=u_{1}^{*}\left(s_{2}\right)$. We then define the trial function
$v(s)=\left\{\begin{array}[]{ll}u_{1}^{*}(s)&\text{for}\,\,\,s\in\left(s_{1},s_{2}\right),\\ z^{*}(s)&\text{for}\,\,\,s\in\left[0,s_{1}\right]\cup\left[s_{2},\left|B_{R}\right|\right].\end{array}\right$.
Then, by reasoning as that of Lemma 3.3, we can still obtain that
$\mu_{1}<\frac{\int_{B_{R}}|\nabla g|^{2}\,\text{d}x+\frac{\sigma}{R}\int_{\partial B_{R}}g^{2}\,\text{d}S}{\int_{B_{R}}g^{2}\,\text{d}x}\leq\mu_{1}.$
This contradiction verifies our desired conclusion.∎
(a) case (i) occurs
(b) case (ii) occurs
Figure 5: The schematic diagrams of Theorem 3.1.
Our result extends the Chiti comparison by incorporating an additional scenario. Although this generalization introduces greater complexity, it nevertheless enables us to derive the desired inequality.
Remark 3.2. If $z^{*}(0)>u_{1}^{*}(0)$, from the argument of Theorem 3.1, we see that the second alternative can be refined as follows: there exists a point $s_{1}\in\left(0,\left|B_{R}\right|\right)$ such that $z^{*}\left(s_{1}\right)=u_{1}^{*}\left(s_{1}\right)$ and
$\left\{\begin{array}[]{ll}z^{*}(s)>u_{1}^{*}(s)&\text{for}\,\,\,\left[0,s_{1}\right),\\ z^{*}(s)<u_{1}^{*}(s)&\text{for}\,\,\,\left(s_{1},\left|B_{R}\right|\right).\end{array}\right$.
4 Proof of Theorem 1.1
Let $f$ and $g$ be two measurable functions defined on $\Omega$. Recall that
$f_{*}(s)=f^{*}(|\Omega|-s)$
is the increasing rearrangement of $f$ [5, 7]. The spherical increasing rearrangement of $f$ is the function
$f_{\star}(x)=f_{*}\left(C_{N}|x|^{N}\right)$
defined on the ball $\Omega^{*}$. From the definition of rearrangements, it follows directly that
$(fg)^{\star}=f^{\star}g^{\star}\,\,\,\text{and}\,\,\,(fg)_{\star}=f_{\star}g_{\star}.$
To prove Theorem 1.1, we first recall the following two properties of spherical rearrangements [6, 7, 26], which will be used in the sequel:
(I) if $f=f(|x|)$ is nonnegative and decreasing (or increasing) as a function of $r=|x|$ for $x\in\Omega$, then $f^{\star}(r)\leq f(r)$ (or $f_{\star}(r)\geq f(r)$) for $r$ between 0 and the radius of $\Omega^{*}$, with strict inequality for $r\in\left(\min_{x\not\in}|x|,r^{*}\right)$ if $f$ is strictly monotone there, where $r^{*}$ denote the radius of $\Omega^{*}$;
(II) if $f$ and $g$ are nonnegative functions, then
$\int_{\Omega^{*}}f_{\star}g^{\star}\,\text{d}x\leq\int_{\Omega}fg\,\text{d}x\leq\int_{\Omega^{*}}f^{\star}g^{\star}\,\text{d}x.$ (4.1)
Furthermore, when $\Omega=\Omega^{*}$, if $f$ is decreasing, then by the definition of $f^{\star}$, we have that
$f^{\star}(r)=f(r)\,\,\,\text{for}\,\,\,r\in\left[0,r^{*}\right].$
Similarly, if $f$ is increasing, then from the definition of $f_{\star}$, it follows that
$f_{\star}(r)=f(r)\,\,\,\text{for}\,\,\,r\in\left[0,r^{*}\right].$
In other words, for a nonnegative radial function $f$ defined on a ball that is monotone with respect to $r=|x|$, the spherical rearrangement coincides with $f$ itself.
Proof of Theorem 1.1. We begin by recalling the Rayleigh quotient for $\mu_{2}$
$\mu_{2}=\inf_{\varphi\in H^{1}(\Omega)\setminus\{0\}}\frac{\int_{\Omega}|\nabla\varphi|^{2}\,\text{d}x+\sigma\int_{\partial\Omega}\varphi^{2}\,\text{d}S}{\int_{\Omega}\varphi^{2}\,\text{d}x},$
where the trial function $\varphi$ is taken orthogonal to the first eigenfunction $u_{1}$ in $L^{2}(\Omega)$.
Let $P=\varphi/u_{1}$. Then $P$ satisfies
$\int_{\Omega}Pu_{1}^{2}\,\text{d}x=0.$
Using the identity $|\nabla\varphi|^{2}=u_{1}^{2}|\nabla P|^{2}+2u_{1}P\nabla P\nabla u_{1}+P^{2}\left|\nabla u_{1}\right|^{2}$, we obtain that
$\mu_{2}=\inf_{\varphi\in H^{1}(\Omega)\setminus\{0\}}\frac{\int_{\Omega}u_{1}^{2}|\nabla P|^{2}\,\text{d}x+2\int_{\Omega}u_{1}P\nabla P\nabla u_{1}\,\text{d}x+\int_{\Omega}P^{2}\left|\nabla u_{1}\right|^{2}\,\text{d}x+\sigma\int_{\partial\Omega}\varphi^{2}\,\text{d}S}{\int_{\Omega}\varphi^{2}\,\text{d}x}.$
Recall that $u_{1}$ satisfies
$\left\{\begin{array}[]{ll}-\Delta u_{1}=\mu_{1}u_{1}&\text{in}\,\,\,\Omega,\\ \partial_{\texttt{n}}u_{1}+\sigma u_{1}=0&\text{on}\,\,\,\partial\Omega.\end{array}\right$.
Multiplying both sides of the equation for $u_{1}$ by $u_{1}P^{2}$ and integrating by parts yields
$2\int_{\Omega}u_{1}P\nabla P\nabla u_{1}\,\text{d}x+\int_{\Omega}P^{2}\left|\nabla u_{1}\right|^{2}\,\text{d}x+\sigma\int_{\partial\Omega}\varphi^{2}\,\text{d}S=\mu_{1}\int_{\Omega}\varphi^{2}\,\text{d}x.$
Substituting this into the previous expression gives
$\mu_{2}-\mu_{1}=\inf_{Pu_{1}\in H^{1}(\Omega)\setminus\{0\}}\frac{\int_{\Omega}u_{1}^{2}|\nabla P|^{2}\,\text{d}x}{\int_{\Omega}P^{2}u_{1}^{2}\,\text{d}x}.$ (4.2)
In the two-dimensional case, this variational characterization was previously established by Payne and Schaefer [39].
Following the approach in [6, 7], we consider $N$ trial functions $P=P_{i}$ for $i=1,\ldots,N$, defined by
$P_{i}=g(r)\frac{x_{i}}{r}$
where $g(r)$ is a positive function of the radial variable $r=|x|$ (to be determined later), and $x_{i}$ denotes the $i$th component of $x$. The function $g$ will be chosen to be continuous, differentiable and bounded on $(0,+\infty)$. By an appropriate choice of the origin, as in [7], we have that
$\int_{\Omega}P_{i}u_{1}^{2}\,\text{d}x=0$
for $i=1,\ldots,N$.
Taking $P=P_{i}$ in (4.2), we obtain that
$\left(\mu_{2}-\mu_{1}\right)\int_{\Omega}P_{i}^{2}u_{1}^{2}\,\text{d}x\leq\int_{\Omega}u_{1}^{2}|\nabla P_{i}|^{2}\,\text{d}x.$
Summing over $i=1,\ldots,N$ and using the identity $\sum_{i=1}^{N}P_{i}^{2}=g^{2}$, it follows that
$\mu_{2}-\mu_{1}\leq\frac{\int_{\Omega}\sum_{i=1}^{N}|\nabla P_{i}|^{2}u_{1}^{2}\,\text{d}x}{\int_{\Omega}g^{2}(r)u_{1}^{2}\,\text{d}x}.$
As shown in [7], we see that
$\sum_{i=1}^{N}|\nabla P_{i}|^{2}=g^{\prime 2}+(N-1)\frac{g^{2}}{r^{2}}.$
Substituting this into the numerator yields the gap inequality
$\mu_{2}-\mu_{1}\leq\frac{\int_{\Omega}\left(g^{\prime 2}+(N-1)\frac{g^{2}}{r^{2}}\right)u_{1}^{2}\,\text{d}x}{\int_{\Omega}g^{2}(r)u_{1}^{2}\,\text{d}x},$ (4.3)
which is of the same form as that derived for the fixed membrane problem in [7].
Define the function
$w(t)=\left\{\begin{array}[]{ll}\frac{J_{\nu+1}(\beta t)}{J_{\nu}(\alpha t)}&\text{for}\,\,\,0\leq t<1,\\ \frac{J_{\nu+1}(\beta)}{J_{\nu}(\alpha)}&\text{for}\,\,\,t\geq 1,\end{array}\right$.
where $\alpha=k_{\nu,1}$ and $\beta=k_{\nu+1,1}$. It is not difficult to verify that $w$ is continuously differentiable on $0,+\infty)$. Let $\gamma=1/R=\sqrt{\mu_{1}}/\alpha$, and choose
$g(r)=w(\gamma r).$
Substituting $g$ into inequality (4.3) yields
$\mu_{2}-\mu_{1}\leq\frac{\mu_{1}}{\alpha^{2}}\frac{\int_{\Omega}B(\gamma r)u_{1}^{2}\,\text{d}x}{\int_{\Omega}w^{2}(\gamma r)u_{1}^{2}\,\text{d}x},$ (4.4)
where
$B(t)=w^{\prime}(t)^{2}+(2\nu+1)\frac{w^{2}(t)}{t^{2}}.$
By Proposition 2.1, the function $w(t)$ is increasing and $B(t)$ is decreasing on $[0,1]$. Remark 2.2 implies that $w(t)$ is positive on $(0,1]$ and $B(t)$ is positive on $[0,1]$. Moreover, by their definitions, these monotonicity properties extend to the entire interval $0,+\infty)$.
Applying inequality (4.1), we obtain that
$\int_{\Omega}B(\gamma r)u_{1}^{2}\,\text{d}x\leq\int_{\Omega^{*}}B^{\star}(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x$
and
$\int_{\Omega}w^{2}(\gamma r)u_{1}^{2}\,\text{d}x\geq\int_{\Omega^{*}}w_{\star}^{2}(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x.$
Furthermore, since the spherical rearrangement of a nonnegative, radial, and monotonous function defined on a ball coincides with the function itself, it follows that
$\int_{\Omega^{*}}B^{\star}(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x=\int_{\Omega^{*}}B(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x$
and
$\int_{\Omega^{*}}w_{\star}^{2}(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x=\int_{\Omega^{*}}w^{2}(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x.$
We further claim that
$\int_{\Omega^{*}}B(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x\leq\int_{B_{R}}B(\gamma r)z^{2}\,\text{d}x$ (4.5)
and
$\int_{\Omega^{*}}w^{2}(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x\geq\int_{B_{R}}w^{2}(\gamma r)z^{2}\,\text{d}x.$ (4.6)
To verify this claim, we first show that
$\int_{\Omega^{*}}f(r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x\geq\int_{B_{R}}f(r)z^{2}\,\text{d}x\,\,\,\text{if $f$ is increasing}$ (4.7)
and the reverse inequality holds if $f$ is decreasing.
By Theorem 3.1, when $\sigma\geq\sigma_{*}$, the functions $u_{1}^{\star}(r)$ and $z(r)$ satisfy either
(i) $z(r)\geq u_{1}^{\star}(r)$ in $\left[0,R\right]$, or
(ii) there exists a point $r_{1}\in\left(0,R\right)$ such that
$\left\{\begin{array}[]{ll}z(r)\geq u_{1}^{\star}(r)&\text{for}\,\,\,\left[0,r_{1}\right],\\ z(r)\leq u_{1}^{\star}(r)&\text{for}\,\,\,\left[r_{1},R\right],\end{array}\right$.
where $z^{*}(s)=z(r)$ with $s=C_{N}|x|^{N}$.
Thus we have that
$\int_{B_{R}}f(r)z^{2}\,\text{d}x-\int_{\Omega^{*}}f(r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x=$ $NC_{N}\left(\int_{0}^{R}f(r)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r\right$.
$\left.-\int_{R}^{r^{*}}f(r)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right).$
If $z(r)\geq u_{1}^{\star}(r)$ in $\left[0,R\right]$, we have that
$\int_{B_{R}}f(r)z^{2}\,\text{d}x-\int_{\Omega^{*}}f(r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x$
$\leq\,$ $NC_{N}\left(\int_{0}^{R}f(R)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}f(R)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $NC_{N}f(R)\left(\int_{0}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $f(R)\left(\int_{B_{R}}z^{2}\,\text{d}x-\int_{\Omega^{*}}\left(u_{1}^{\star}\right)^{2}\,\text{d}x\right)$
$=$ $0.$
If there exists a point $r_{1}\in\left(0,R\right)$ such that
$\left\{\begin{array}[]{ll}z(r)\geq u_{1}^{\star}(r)&\text{for}\,\,\,\left[0,r_{1}\right],\\ z(r)\leq u_{1}^{\star}(r)&\text{for}\,\,\,\left[r_{1},R\right],\end{array}\right$.
then we have that
$\int_{B_{R}}f(r)z^{2}\,\text{d}x-\int_{\Omega^{*}}f(r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x$
$=$ $NC_{N}\left(\int_{0}^{r_{1}}f(r)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r+\int_{r_{1}}^{R}f(r)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r\right$.
$\left.-\int_{R}^{r^{*}}f(r)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$\leq\,$ $NC_{N}\left(\int_{0}^{r_{1}}f\left(r_{1}\right)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r+f\left(r_{1}\right)\int_{r_{1}}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r\right$.
$\left.-f\left(r_{1}\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $NC_{N}f\left(r_{1}\right)\left(\int_{0}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $f\left(r_{1}\right)\left(\int_{B_{R}}z^{2}\,\text{d}x-\int_{\Omega^{*}}\left(u_{1}^{\star}\right)^{2}\,\text{d}x\right)$
$=$ $0.$
Thus, inequality (4.7) is established. The corresponding result for decreasing $f$ follows by a similar argument. Applying (4.7) and its counterpart for decreasing $f$, we confirm the desired inequalities (4.5)–(4.6).
Therefore, we conclude that
$\int_{\Omega}B(\gamma r)u_{1}^{2}\,\text{d}x\leq\int_{B_{R}}B(\gamma r)z^{2}\,\text{d}x$
and
$\int_{\Omega}w^{2}(\gamma r)u_{1}^{2}\,\text{d}x\geq\int_{B_{R}}w^{2}(\gamma r)z^{2}\,\text{d}x.$
Combining these with (4.4), we obtain that
$\mu_{2}-\mu_{1}\leq$ $\frac{\mu_{1}}{\alpha^{2}}\frac{\int_{B_{R}}B(\gamma r)z^{2}\,\text{d}x}{\int_{B_{R}}w^{2}(\gamma r)z^{2}\,\text{d}x}$
$=$ $\frac{\mu_{1}}{\alpha^{2}}\frac{\int_{0}^{R}B(\gamma r)J_{\nu}^{2}\left(\sqrt{\mu_{1}}r\right)r\,\text{d}r}{\int_{0}^{R}w^{2}(\gamma r)J_{\nu}^{2}\left(\sqrt{\mu_{1}}r\right)r\,\text{d}r}$
$=$ $\frac{\mu_{1}}{\alpha^{2}}\frac{\int_{0}^{1}B(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r}{\int_{0}^{1}w^{2}(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r}$
where the last equality follows from a change of variables and the relation $\gamma R=1$.
Using the antidifferentiation formula of [7, Formula B.3], we obtain that
$\frac{\mu_{1}}{\alpha^{2}}\frac{\int_{0}^{1}B(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r}{\int_{0}^{1}w^{2}(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r}=\frac{\mu_{1}}{\alpha^{2}}\frac{{rA(r)J_{\nu+1}^{2}(\beta r)}\big|_{0}^{1}+\left(\beta^{2}-\alpha^{2}\right)\int_{0}^{1}J_{\nu+1}^{2}(\beta r)r\,\text{d}r}{\int_{0}^{1}w^{2}(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r},$
where $A(r)=(\ln w(r))^{\prime}$. Since $J_{\nu+1}(0)=0$, $A(1)=q(1)=0$ and $w(r)=J_{\nu+1}(\beta r)/J_{\nu}(\alpha r)$, we find that
$\frac{\mu_{1}}{\alpha^{2}}\frac{\int_{0}^{1}B(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r}{\int_{0}^{1}w^{2}(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r}=\frac{\mu_{1}}{\alpha^{2}}\left(\beta^{2}-\alpha^{2}\right)\frac{\int_{0}^{1}J_{\nu+1}^{2}(\beta r)r\,\text{d}r}{\int_{0}^{1}w^{2}(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r}=\frac{\mu_{1}}{\alpha^{2}}\left(\beta^{2}-\alpha^{2}\right).$
Thus we have that
$\mu_{2}-\mu_{1}\leq\frac{\mu_{1}}{\alpha^{2}}\left(\beta^{2}-\alpha^{2}\right),$
which implies that
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{\beta^{2}}{\alpha^{2}}=\frac{k_{\nu+1}^{2}}{k_{\nu}^{2}}.$
This completes the proof of the desired inequality.
We now prove that equality holds only when the domain $\Omega$ is a ball. Since $\left|B_{R}\right|<|\Omega|$, it follows that $R<r^{*}$. We claim that
$\int_{\Omega^{*}}f(r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x>\int_{B_{R}}f(r)z^{2}\,\text{d}x\,\,\,\text{if $f$ is strictly increasing on $[0,R]$}$ (4.8)
and the reverse inequality holds if $f$ is strictly decreasing on $[0,R]$, where $f$ is extended to $R,r^{*}]$ as the constant $f(R)$ with $f(R)>0$. Indeed, if $z(r)\geq u_{1}^{\star}(r)$ in $\left[0,R\right]$, there are two case: either $z(r)\equiv u_{1}^{\star}(r)$ in $\left[0,R\right]$ or $z(r)\geq u_{1}^{\star}(r)$ and $z(r)\not\equiv u_{1}^{\star}(r)$ in $\left[0,R\right]$. In fact, the former case is impossible. Otherwise, we find that
$\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s<\int_{0}^{M}\left(u_{1}^{*}\right)^{2}(s)\,\text{d}s=\int_{0}^{\left|B_{R}\right|}\left(z^{*}\right)^{2}(s)\,\text{d}s,$
which is a contradiction. Hence, we must have that $z(r)\geq u_{1}^{\star}(r)$ and $z(r)\not\equiv u_{1}^{\star}(r)$ in $\left[0,R\right]$. In this case, one has that
$\int_{B_{R}}f(r)z^{2}\,\text{d}x-\int_{\Omega^{*}}f(r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x$
$=$ $NC_{N}\left(\int_{0}^{R}f(r)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-f(R)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right).$
Since $f$ is strictly increasing, we find that
$\int_{0}^{R}(f(R)-f(r))\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r>0.$
It then follows that
$\int_{B_{R}}f(r)z^{2}\,\text{d}x-\int_{\Omega^{*}}f(r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x$
$<\,$ $NC_{N}\left(\int_{0}^{R}f(R)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}f(R)\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $NC_{N}f(R)\left(\int_{0}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r-\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $f(R)\left(\int_{B_{R}}z^{2}\,\text{d}x-\int_{\Omega^{*}}\left(u_{1}^{\star}\right)^{2}\,\text{d}x\right)$
$=$ $0.$
If there exists a point $r_{1}\in\left(0,R\right)$ such that
$\left\{\begin{array}[]{ll}z(r)\geq u_{1}^{\star}(r)&\text{for}\,\,\,\left[0,r_{1}\right],\\ z(r)\leq u_{1}^{\star}(r)&\text{for}\,\,\,\left[r_{1},R\right],\end{array}\right$.
then we obtain that
$\int_{B_{R}}f(r)z^{2}\,\text{d}x-\int_{\Omega^{*}}f(r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x$
$\leq\,$ $NC_{N}\left(\int_{0}^{r_{1}}f\left(r_{1}\right)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r+f\left(r_{1}\right)\int_{r_{1}}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r\right$.
$\left.-f\left(R\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right).$
Since $f$ is strictly increasing, $r_{1}<R<r^{*}$ and $u_{1}^{\star}>0$, it follows that
$f\left(R\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r>f\left(r_{1}\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r.$
Thus, we have that
$\int_{B_{R}}f(r)z^{2}\,\text{d}x-\int_{\Omega^{*}}f(r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x$
$<\,$ $NC_{N}\left(\int_{0}^{r_{1}}f\left(r_{1}\right)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r+f\left(r_{1}\right)\int_{r_{1}}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r\right$.
$\left.-f\left(r_{1}\right)\int_{R}^{r^{*}}\left(u_{1}^{\star}\right)^{2}r^{N-1}\,\text{d}r\right)$
$=$ $f\left(r_{1}\right)\left(\int_{B_{R}}z^{2}\,\text{d}x-\int_{\Omega^{*}}\left(u_{1}^{\star}\right)^{2}\,\text{d}x\right)$
$=$ $0.$
We therefore conclude that the desired inequality (4.8) holds. The reversed inequality for the case when $f$ is strictly decreasing follows by an analogous argument.
By Proposition 2.2, $w(t)$ is strictly increasing and $B(t)$ is strictly decreasing on $[0,1]$. Leveraging inequality (4.8) and the reversed form for decreasing functions, we find that
$\int_{\Omega^{*}}B(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x<\int_{B_{R}}B(\gamma r)z^{2}\,\text{d}x$
and
$\int_{\Omega^{*}}w^{2}(\gamma r)\left(u_{1}^{\star}\right)^{2}\,\text{d}x>\int_{B_{R}}w^{2}(\gamma r)z^{2}\,\text{d}x.$
This implies that
$\frac{\mu_{2}}{\mu_{1}}<\frac{\beta^{2}}{\alpha^{2}}=\frac{k_{\nu+1}^{2}}{k_{\nu}^{2}}.$
In conclusion, we have that
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{k_{\nu+1}^{2}}{k_{\nu}^{2}},$
with equality if and only if $\Omega=B_{R}$.∎
Remark 4.1. When $\Omega$ is a unit ball, taking $g=w(\gamma r)$ with $w(r)=J_{\nu+1}(\beta r)/J_{\nu}(\alpha r)$ for $0\leq r\leq 1$, (4.3) reduces to
$\beta^{2}-\alpha^{2}=\frac{\int_{0}^{1}B(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r}{\int_{0}^{1}w^{2}(r)J_{\nu}^{2}\left(\alpha r\right)r\,\text{d}r}.$
This also implies
$\frac{\mu_{2}}{\mu_{1}}\leq\frac{k_{\nu+1}^{2}}{k_{\nu}^{2}}.$
Moreover, the proof of the “only if” part can be adapted from the arguments in [6, 7].
Remark 4.2. The assumptions on $f$ in (4.8) and its reversed form can be weakened to: $f$ is strictly monotonic on some subinterval with non-zero length of $\left[r_{1},R\right]$. Indeed, in this case, when $f$ is increasing, we still have that
$\int_{0}^{R}(f(R)-f(r))\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r>0.$
Moreover, in view of Remarks 3.1–3.2, when $f$ is increasing, one has that
$\int_{0}^{r_{1}}\left(f\left(r_{1}\right)-f(r)\right)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r$
is nonnegative and
$\int_{r_{1}}^{R}\left(f\left(r_{1}\right)-f(r)\right)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r$
is positive. This follows that
$\int_{0}^{r_{1}}f\left(r\right)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r+\int_{r_{1}}^{R}f\left(r\right)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r$
$<\int_{0}^{r_{1}}f\left(r_{1}\right)\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r+f\left(r_{1}\right)\int_{r_{1}}^{R}\left(z^{2}-\left(u_{1}^{\star}\right)^{2}\right)r^{N-1}\,\text{d}r.$
The reversed form is similar.
5 Proof of Theorem 1.2
Combining equations (2.8) and (2.9), we derive an expression for the derivative of $\beta/\alpha$. Applying Lemma 2.1 to this result then yields the desired conclusion.
Proof of Theorem 1.2. Define
$F(\sigma,\alpha)=\frac{\alpha J_{\nu+1}(\alpha)}{J_{\nu}(\alpha)}-\sigma$
and
$G(\sigma,\beta)=\frac{\beta J_{\nu+2}(\beta)}{J_{\nu+1}(\beta)}-\sigma-1.$
By (2.8) and (2.9), we see that $F(\sigma,\alpha)=0$ and $G(\sigma,\beta)=0$. It is easily verified that both $F$ and $G$ are differentiable with continuous partial derivatives, and that $F_{\sigma}(\sigma,\alpha)=-1$ and $G_{\sigma}(\sigma,\beta)=-1$. Therefore, by the implicit function theorem, $\alpha$ and $\beta$ are continuously differentiable with respect to $\sigma$. Moreover, differentiating the identities $F(\sigma,\alpha)=0$ and $G(\sigma,\beta)=0$ yields
$\alpha^{\prime}(\sigma)=-\frac{F_{\sigma}}{F_{\alpha}}=\frac{1}{\frac{J_{\nu+1}(\alpha)}{J_{\nu}(\alpha)}+\alpha\frac{J_{\nu+1}^{\prime}(\alpha)J_{\nu}(\alpha)-J_{\nu+1}(\alpha)J_{\nu}^{\prime}(\alpha)}{J_{\nu}^{2}(\alpha)}}$
and
$\beta^{\prime}(\sigma)=-\frac{G_{\sigma}}{G_{\beta}}=\frac{1}{\frac{J_{\nu+2}(\beta)}{J_{\nu+1}(\beta)}+\beta\frac{J_{\nu+2}^{\prime}(\beta)J_{\nu+1}(\beta)-J_{\nu+2}(\beta)J_{\nu+1}^{\prime}(\beta)}{J_{\nu+1}^{2}(\beta)}}.$
Using (2.5) and the identity
$J_{\nu+1}^{\prime}(z)=J_{\nu}(z)-\frac{\nu+1}{z}J_{\nu+1}(z),$
we further obtain that
$\alpha^{\prime}(\sigma)=$ $\frac{1}{\frac{J_{\nu+1}(\alpha)}{J_{\nu}(\alpha)}+\alpha\frac{\left(J_{\nu}(\alpha)-\frac{\nu+1}{\alpha}J_{\nu+1}(\alpha)\right)J_{\nu}(\alpha)-J_{\nu+1}(\alpha)\left(-J_{\nu+1}(\alpha)+\frac{\nu}{\alpha}J_{\nu}(\alpha)\right)}{J_{\nu}^{2}(\alpha)}}$
$=$ $\frac{1}{\frac{J_{\nu+1}(\alpha)}{J_{\nu}(\alpha)}-(2\nu+1)\frac{J_{\nu+1}(\alpha)}{J_{\nu}(\alpha)}+\alpha\frac{J_{\nu+1}^{2}(\alpha)}{J_{\nu}^{2}(\alpha)}+\alpha}$
$=$ $\frac{1}{\alpha\frac{J_{\nu+1}^{2}(\alpha)}{J_{\nu}^{2}(\alpha)}-2\nu\frac{J_{\nu+1}(\alpha)}{J_{\nu}(\alpha)}+\alpha}$
and
$\beta^{\prime}(\sigma)=$ $\frac{1}{\frac{J_{\nu+2}(\beta)}{J_{\nu+1}(\beta)}+\beta\frac{J_{\nu+2}^{\prime}(\beta)J_{\nu+1}(\beta)-J_{\nu+2}(\beta)J_{\nu+1}^{\prime}(\beta)}{J_{\nu+1}^{2}(\beta)}}$
$=$ $\frac{1}{\frac{J_{\nu+2}(\beta)}{J_{\nu+1}(\beta)}+\beta\frac{\left(J_{\nu+1}(\beta)-\frac{\nu+2}{\beta}J_{\nu+2}(\beta)\right)J_{\nu+1}(\beta)-J_{\nu+2}(\beta)\left(-J_{\nu+2}(\beta)+\frac{\nu+1}{\beta}J_{\nu+1}(\beta)\right)}{J_{\nu+1}^{2}(\beta)}}$
$=$ $\frac{1}{\frac{J_{\nu+2}(\beta)}{J_{\nu+1}(\beta)}-(2\nu+3)\frac{J_{\nu+2}(\beta)}{J_{\nu+1}(\beta)}+\beta\frac{J_{\nu+2}^{2}(\beta)}{J_{\nu+1}^{2}(\beta)}+\beta}$
$=$ $\frac{1}{\beta\frac{J_{\nu+2}^{2}(\beta)}{J_{\nu+1}^{2}(\beta)}-2(\nu+1)\frac{J_{\nu+2}(\beta)}{J_{\nu+1}(\beta)}+\beta}.$
Using (2.8) and (2.9) again we obtain that
$\alpha^{\prime}(\sigma)=$ $\frac{1}{\frac{\sigma^{2}}{\alpha}-2\nu\frac{\sigma}{\alpha}+\alpha}=\frac{\alpha}{\sigma^{2}-2\nu\sigma+\alpha^{2}}$
and
$\beta^{\prime}(\sigma)=$ $\frac{1}{\frac{(\sigma+1)^{2}}{\beta}-2(\nu+1)\frac{\sigma+1}{\beta}+\beta}$
$=$ $\frac{\beta}{(\sigma+1)^{2}-2(\nu+1)(\sigma+1)+\beta^{2}}$
$=$ $\frac{\beta}{\sigma^{2}-2\nu\sigma-2\nu-1+\beta^{2}}.$
From [17] or [22], it is known that both $\alpha$ and $\beta$ are strictly increasing in $\sigma$. Using the identity $\alpha^{2}\left(\beta/\alpha\right)^{\prime}(\sigma)=\beta^{\prime}\alpha-\alpha^{\prime}\beta$, we obtain that
$\frac{\alpha^{2}\left(\beta/\alpha\right)^{\prime}(\sigma)}{\alpha^{\prime}(\sigma)\beta^{\prime}(\sigma)}=$ $\frac{\alpha}{\alpha^{\prime}}-\frac{\beta}{\beta^{\prime}}$
$=$ $\sigma^{2}-2\nu\sigma+\alpha^{2}-\left(\sigma^{2}-2\nu\sigma-2\nu-1+\beta^{2}\right)$
$=$ $2\nu+1+\alpha^{2}-\beta^{2}<0$
via Lemma 2.1. Hence, we conclude that $\beta/\alpha$ is strictly decreasing in $\sigma$.
As in [17], define the function
$h(k)=\frac{kJ_{\nu+2}(k)}{J_{\nu+1}(k)}-1.$
From [17], it is known that $h$ is strictly increasing on $\left[0,j_{\nu+1,1}\right)$ with h(0)=1h(0)=-1. Therefore, there exists a unique zero $k_{*}\in\left(0,j_{\nu+1,1}\right)$. Following the argument in [17, Theorem 1.1], we obtain that
$\lim_{\sigma\rightarrow 0}\alpha=0\,\,\,\text{and}\,\,\,\lim_{\sigma\rightarrow 0}\beta=k_{*},$
which implies
$\lim_{\sigma\rightarrow 0}\frac{\beta}{\alpha}=+\infty.$
Moreover, since
$\lim_{\sigma\rightarrow+\infty}\frac{\beta}{\alpha}=\frac{j_{\nu+1,1}}{j_{\nu,1}},$
we conclude that $\beta/\alpha$ is strictly decreasing from $\infty$ to $j_{\nu+1,1}/j_{\nu,1}$ as $\sigma$ increases from 0 to $\infty$. This establishes the desired result. ∎
In the above proof, relations (2.8)–(2.9) and Lemma 2.1 play essential roles. It is also worth noting that the argument used in Theorem 1.2 remains valid for the case $N=1$.
The conflicts of interest statement and Data Availability statement.
There is not any conflict of interest. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Acknowledgment
The authors express their sincere gratitude to Professor Richard Laugesen for his invaluable guidance and insightful discussions regarding this work. They are especially thankful to him for identifying issues and offering revision suggestions on the initial draft, as well as for providing a wealth of relevant literature and materials.